A Complete Bibliography of the Journal of Econometrics (1970–1979)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.01

Title word cross-reference

$10.00 [Tak73, Maz75]. $13.20 [O’B75]. $18.75 [Lee73, Pag73a]. $19.00 [Blo73]. $22.50 [Bal75]. $30.00 [Dag77, Wil74]. $31.50 [Ric75b]. $32.50 [Tag75]. $41.50 [Aig73b]. $72.00 [Aig75]. r̄^{2} [McE77a]. K [Sch76, FS74, Mar75]. R^{2} [GN76]. R_{y.x}^{2} [McE77a].

* [Wal77].

-class [FS74, Mar75]. -matrix-class [Sch76].

1973 [Ano73f, Ano73d, Ano73c, Ano73e]. 1974 [Ano74h, Ano74f, Ano74e, Ano74g]. 1975 [Ano75i, Ano75g, Ano75h, Ano75j]. 1976 [Ano76g, Ano76e, Ano76f, Ano76h]. 1977 [Ano77i, Ano77j, Ano77k, Ano77n, Ano77m, Ano77l]. 1978 [Ano78l, Ano78h, Ano78m, Ano78i, Ano78k, Ano78j]. 1979 [Ano79o, Ano79u, Ano79t, Ano79q, Ano79r, Ano79s, Ano79p].

2.30 [Poi75a]. 243 [Mit77]. 2nd [O’H77].

91 [Tri75].

ability [Cha77]. Abrahamse [Neu77]. Acknowledgement [Ano73a, Ano74a, Ano75a, Ano76a, Ano77a, Ano79a]. ad [Gil75]. adaptive [DG78]. added [DM77]. additive [Koe79]. adjustment [DG78, Hat74]. adjustment-adaptive [DG78]. adjustments [Cra79]. administrative [Bal75]. advertising [Blo73]. after [BJY73, JYB73, TW75]. aggregated [Far79, Hsi79]. aggregates [Ros77]. aggregation [Bre73, Den78, Lov73, Sas78, Wei78]. aggregative [Ho79]. algebraic [FS74]. allocation [HS76]. allocative [SL79]. Almon [HH78, ZW73]. Alternative [Gui74, FG78]. alternatives [HH78]. Amsterdam [Aig73a, Aig73b, Aig75, Bal75, Blo73, Cho74, Cho77, Die75, Hur73, Lee73, Mad76, Mit77, O’H77, Pag73a, Ric75b, Tag75, Tak73, Wil74]. analysis [Aig75, Aig79a, Bal75, Blo73, (Mu77, Dag73, Drè77, GV77, GERA79, HC77, HR77, Hun79, Laf77, LT78b, Nef79, Neu77, Plo79, Sas78, ZW73, ZP74, Cho77, Jud75, Tag75, Wil74]. Angeles [MMA79]. Announcement [Ano78a, Ano79b]. annual [Gil77]. Anticipated [Aig79c]. Aoki [Cho77]. appearance [GM79b]. application [AGS79, And74, Att77, Cho74, Had76, Kak77, Ken75, LT78a, MP78, Tsu77]. applications [Bal75, Mai78, Tak73]. applied [Hur73, Tsu76, Jud75]. approach [BKK79, Dag73, Gro75, HKP79, PP79, SW78, Uri79, Wal77, Zel78]. appropriate [GN76]. approx [Dag77]. approximation [Men77, Phi77b]. approximations [Rob77]. April [Ano78l, Ano79o]. Arima [Ros77]. ARMA [Bre73, But76, Rei79]. ARMAX [Bre73]. Articles [Ano75b]. Arts [Gav76]. aspects [TTC77, Die75]. assessment [GR74]. associated [And75]. Asymptotic [DTH78, DE74, GB78, Men77]. asymptotics [Mar75]. attributes [WG78]. August [Ano75i, Ano76g, Ano78h, Ano79u]. Australian [HF73]. Autocorrelated [(Mu77, Cor79, Dei78, GM79b, Hen79]. Autocorrelation [Ken75, DLA78, Fit73, GK78, LT75, SW78]. autoregression [Phi79b]. autoregressive [BM78, DM78, DE74, Fit73, Gil75, Gui74, Hat74, Hat76, Men77, Nel76, Saw78, Spe79]. autoregressive-moving [Nel76]. average [ASW77, DM78, Fit73, LT78b, Men77, Nel74, Nel76, NS79, PS77].

Ball [Ric75b]. band [Hyl77]. bank [Aig73c, Lap76]. based [McC74, NS79]. Bayes [GR79a, MS78, SR75, SR78]. Bayesian [Mad76, Aig79a, Drè77, FG78, Gil75, GR74, GDP79, Gro75, Gut76, HV77, SM75, TT76, Tsu76, Tsu77, ZW73]. behavior [App79, Sas78, Lap76]. behaviour [Aig73c, HH74, Hen79, Tri73]. Berkeley [Sta74]. Best [Bal73, Aig79a]. between [Gil75]. bias [Lov73, NS78, Spe75]. bibliography [Sow73]. bilinear [Poi75b]. binary [Aig73d, Lan79, Wes74]. bivariate [TTC77]. blue [Sch75b]. board [Ano76b, Ano78c, Pan76, Ano74b, Ano75c, Ano77b, Ano77c, Ano78d, Ano79e, Ano79f, Ano79g]. Book [Aig73a, Aig73b, Aig75, Bal75, Blo73, Cho74, Cho77, Dag77, Die75, Gav76, Hur73, Jud75, Kad76, Lap76, Lee73, Lev77a, Mad76, Maz75, Mit77, O’B75, O’H77, Pag73a, Poi75a, Ric75b, Sch75a, Sta74, Tag75, Tak73, Tri75, Wei75, Wil74]. both [Hsi79]. bounded [Lev77b]. Box [And75, NG79]. brief [Aig79b, SD79]. business [Wei75].

C [Lev77a]. California [Sta74]. call [Pan76]. Canadian [Wei75, DM77, Fus77]. Canonical [Vin76]. Carlo [DM78, Gil75, HH74, Sow73, YJ76]. case [Dei78, PS77, Sas78, Sch76, WG78]. Causality [PH77, PH79, Pri79]. Censored [Nel77]. censoring [Nel77]. CES [HV77, TT76]. change [Wil79]. changes [Pan76, UM78]. characteristics [DLA78]. Characterization [PH77, GM79a, PH79, Pri79]. Chicago [Pan76]. Chilean [Wil74]. choice [KK74, Lan79, Man75, Wes74, WG78]. choosing [FG78]. class [DN73, FS74, Kel75, Mar75, Sch76]. classified [Sow73]. clothing [MP78]. Co [Aig73a, Aig73b, Blo73, Hur73, Lee73, Mad76, Pag73a, Tak73, Poi75a]. coefficient [AGS79, DTH78, GDP79, NN78, PZ78, Vis78]. Coefficients [CN77, MS78, SR75, SR78, SM79, Ull74, Zel78]. combined [Saw73]. comment [Kad78a, PH79]. comments [BJY73]. common [Pan76, SM79]. commuting [Wal78]. comparative [Hyl77, Kle79b]. comparison [Gil75, HP74, Saw73, YJ76]. comparisons [FHM75]. competitiveness [Bes79]. complete [Gew78, Kle79b]. components [Ber79, Had76, Tau79]. composite [Pag73b]. computation [Cor79, Den76]. computational [Bes79]. compute [Sal76]. concentration [Mar75, Phi79a]. conditional [GM79a]. Conditions [Ber79, TZ79]. confidence [Sal76]. conjugate [GR79a]. connected [DTH78]. Connecticut [HKM79]. connection [Pag78]. consequences [Bet77, Bre73]. considerations [Aok74]. Consistency [Gab78, Koh78]. constraints [TTC77]. construction [Rob77]. consumer [CM77, Koe79]. Consumption [HKP77, HS76, Sta74, ZW73]. containing [Att77]. context [GJ79, Tau79]. continuous [Phi73, Rob77]. Contributions [Blo73, Aig75]. control [Aig79a, Aok74, Cho74, Cho77]. correction [Pri79]. correlated [SR78]. correlation [CN77, DTH78, NN78, PZ78, Spe75, Szr78]. Correlations [Gre74]. Corrigenda [Ano78b, Ano79c, Ano79d]. cost [SL79]. courses [Poi75a]. covariance [Bal73, Mag78]. Cowless [Maz75]. Cox [NG79]. criteria [McE77b]. critique [Sta74]. cross [BKK79, Dei78]. cross-equation [Dei78]. cross-section [BKK79]. crossed [FB74]. crossed-error [FB74]. crude [Tsu76]. Cuddy [O’B75]. curves [GERA79]. cycle [Sta74, Tsu76].

D [Kad76, Lap76, Mit77, O’B75, Sch75a]. data [Afr74, But76, CM79, Far79, Gil77, Gre74, HR74, HR77, Ho79, Hsi79, Kak77, TW79]. day [Aig79c, Atk79, GERA79, HKM79, LB79, PW79, Tay79]. December [Ano79s, Ano73f, Ano74h, Ano78m]. decision [Aig79a]. decomposition [Bal75]. demand [Dan79, Fus77, HKM79, HKP79, Kle79b, LB79, LT78a, MP78, MNS78, PW79, Spi77, Tay79, WG78]. demands [SB79]. densities [Drè77]. dependent [Gun74, LT78a, Spe75]. describing [DN73]. Design [Con79, MMA79, Aig79b, Aig79c, Mor79, Pap78, Tag75]. designs [CW79]. determinants [PW79]. determination [Ken75]. determine [GN76]. Determining [De 78]. development [DN73, Wil74]. Dfl [Aig73a, Hur73, O’B75]. dichotomous [Gun74]. difference [Sti76]. Differencing [GR77]. different [MNS78]. direct [Aig79a]. disaggregate [WG78]. disaggregated [Hsi79]. discrete [Rob77]. Discriminating [Gil75]. Discrimination [AR75]. Disequilibrium [LM79, GQ75]. distributed [Cor77, Gut76, SG76, Wei78]. distribution [GB78, GM79a, KvD78, Phi77b, Phi79b, PZ78, Ric75a, Ull74, UM78]. distributions [GM79a]. disturbance [Mag78, Pag73b]. disturbances [(Mu77, GM79b, Hat76, Kle79a, McE77b, Neu77, Rei79, Spe79]. doctoral [Ano79j]. dominance [Aig74]. Donald [Lap76]. Droz [Dag77]. Duality [Bur75]. dummy [Sal76]. Duncan [Kad76]. Durbin [FG78, Spe75]. Dutch [HV77]. dynamic [Bet77, Cho77, DN73, De 78, DE74, Gew78, GB78, Hat74, Hat76, Hen79, Ken75, LM79, MP78, MNS78, Pie75, Rei79, Spe79, SR75, SR78].

earnings [Kie78]. East [Wei75]. Eckhaus [Wil74]. Econometric [SB79, Aig73c, De 78, GERA79, Gro75, Gus78, HKP79, HF73, Ho79, Mai78, Nef79, Pal77, RS77, Uri79, ZP74, Ano78g, Mit77, Pag73a]. Econometrics [Mad76, GN74, LM79, Pap78, Rob77, Sow73, Vin76, Lee73, Ano79a, Gav76, Lev77a]. economic [Aig75, Blo73, Cho74, Cho77, DLA78, Gas75, Gav76, Hat78, JP78, NG79, O’H77, Plo79, Ric75b, Sch75a]. economics [Ano79j, Bal75, Blo73, FH79, Poi75b, Sch75a, Wei75, O’B75]. economy [HF73, Wei75, Wil74]. ed [Mit77, O’H77, Ric75b]. Editorial [ADZ73, AZ78, Ano76b, Ano78c, Ano74b, Ano75c, Ano77b, Ano77c, Ano78d, Ano79e, Ano79f, Ano79g]. Editors [AM79, Ano79h]. eds [Aig73b, Kad76, Pag73a, Wil74]. Education [Cha77]. educational [CMP77]. Edwin [Hur73]. effect [Wei78]. effects [SG76]. efficiencies [SR75, SR78]. efficiency [LT78b, Nel76, Sri73]. Efficient [KvD78, Pag73b, Dhr74, Hat74, Hat76, Hat78, Maz75]. elasticities [Kak77, Kak78]. elasticity [SJ75]. electricity [Aig79c, Atk79, HKM79, HKP77, HKP79, LB79, MMA79, PW79, SB79, Tay79]. ellipsoid [Phi79a]. Elsevier [Aig73a]. Empirical [KK74, Atk79, Gus78, LT78b, Sas78, Wal77, Wal78]. employment [Kie78, UM78]. endogeneity [Fro74]. endogenous [WG78]. energy [Fus77, Mit77]. Engel [Kak77, Kak78]. equation [BKK79, CN77, Dei78, Den78, DE74, DS78b, Ger76, Gew78, GR74, Hau77, Kad76, Kni77, MS78, MV74, Pal77, SR78, ZP74]. equations [Bes79, Con79, Dag76, Den76, Den78, Gal77, GJ79, Hat76, Hen76, Kel75, Rei79, Spi77, Sri73, SD79, Sti76, SM79, Woo79]. equipment [BC73]. Erratum [Ano75d, Ano77d, Ano79i, BF75]. error [BM78, Ber79, Bet77, Fit73, FB74, Ger76, GR74, Gui74, Hol77, NS78, Saw73, YJ76]. Errors [Hau77, Aig73d, Aig74, Cor79, Dei78, GR79a, Hat74, Hen79, Lev77b, Sal76, SJ75, SR78]. errors-of-observation [Aig74]. Essays [Gav76]. establishment [CM79]. estimate [Vis78]. estimated [GM79b]. estimates [Ber79, Cor79, Lea75, Lev77b, McC74]. Estimating [CM77, SL79, CW79, Kak78, Sri73]. Estimation [Att77, Far79, FB74, GQ75, HR77, Hur79, LT78a, MNS78, Men77, PS77, SW78, Sch77, SJ75, Spe79, SD79, SM79, TW75, Zel78, Aig73c, ALS77, Ame78, BM78, Bes79, Bet77, BL75, BJY73, Den76, Fis76, Fus77, Gal77, GJ79, Gas75, GDP79, Gut76, HV77, HR74, Hat78, Hun79, Kad78a, Kak77, KvD78, KK74, Mag78, Man75, Mar76, Maz75, MP78, Nel74, Nel76, Pag73b, Poi75a, Rei79, Rob77, Sch78, Sch75b, SB79, SM75, TT76, Wei78, Woo79]. estimator [Ame74, Ame75, Ame77a, Dhr74, DG78, FS74, Hat74, Neu77, Phi77b, Saw73, Saw78, Sch76]. estimators [Bal73, Cor77, Den78, DE74, GR79a, GS79, Hat76, Hen76, Hen79, Hyl77, JYB73, Kel75, Kni77, Mar75, MS78, Phi77a, Sch76, SR75, SR78, Ull74, YJ76]. European [Ano78g, Ano79j]. Evaluation [HKM79]. evidence [CM79]. Exact [Die76, Saw78]. example [Fus77]. existence [Ber79, DM77, Kni77, MS78]. exogeneity [Gew78]. expansion [Afr74]. Expectations [Cra79, DG78, Gro75, Hun79]. expected [Zel78]. Experience [NG79]. experiment [Aig79c, MMA79]. experimental [Aig79b, CW79, Mor79, Pap78]. experiments [Aig79a, Aig79c, FH79, NMM^{+}79]. extraneous [BW73].

family [Gas75]. February [Ano75g, Ano76e, Ano78i, Ano79t]. Federal [ZW73]. Federally [Kie78]. Fienberg [Mad76]. FIML [Rei79]. final [De 78]. finance [Aig73b]. finite [Hyl77, Men77, Mor79, Pag78, Phi73, Phi77b]. Firm [GR79b]. firms [Sas78]. First [Atk79, Mon78, Gui74, Men77, Nel74, Phi79b]. First-order [Mon78, Gui74, Men77, Nel74, Phi79b]. fiscal [McC74]. fit [Bus79, McE77a, NS79]. five [DLA78]. fixed [PZ78]. flexibility [Wal77]. flexible [BPR77, CM77, Wal77]. food [HS76]. Forecasting [Pie75, Ros77, NG79, Uri79]. forecasts [NS78, Phi79b]. form [De 78, Kni77, NS78, SW78]. forms [BPR77, Gil75, Spi77, Wal77]. Formulation [ALS77]. Foundation [Maz75]. four [LT75]. Fourth [GK78]. Fourth-order [GK78]. Fox [O’H77]. frontier [ALS77, LT78b]. frontiers [SL79]. Full [BM78, Bes79, DE74, Sch76]. full-information [Bes79, Sch76]. function [ALS77, BC73, BL75, CM77, CM79, DM77, HV77, LT78b, MNS78, ZW73]. functional [BPR77, HC77, SW78, Spi77, Wal77]. functions [Cor77, Die75, Has76, JP78, Kle79b, MP78, TT76, Zel78]. Further [GK78, HR77].

G [Aig73b]. Gains [Nel76]. gasoline [MNS78]. general [Ame77a, DTH78, HP74]. generalization [Bus79]. Generalized [Szr78, Dag73, Sch75b, Spi77]. generation [NMM^{+}79]. Geneva [Dag77, Ano78g]. George [Aig75]. Glahn [McE77a]. global [Koh78]. GLS [Mag78]. Goldberger [Kad76]. Goldfeld [Lee73]. Goodness [Bus79, McE77a, NS79]. Goodness-of-fit [Bus79, NS79]. Granger [And75]. graph [FS74]. Griffin [Poi75a]. grouped [Att77, Kak77, Poi75a].

H [Tag75, Wei75]. Haitovosky [Poi75a]. Harmonic [HH78]. Haven [Lap76]. health [Mor79, NMM^{+}79]. Henri [Bal75]. Hester [Lap76]. heteroscedastic [Ame77b]. heteroscedasticity [Hur79]. heteroskedasticity [God78, HP74]. Hilbert [Cor77]. hoc [Gil75]. Holland [Aig73b, Aig75, Bal75, Blo73, Cho74, Cho77, Die75, Hur73, Lee73, Mad76, Mit77, O’H77, Pag73a, Ric75b, Tag75, Tak73, Wil74]. Holt [Jud75]. homogeneity [TW75]. honor [Mad76]. honour [Gav76]. Hooper [McE77a]. household [HS76]. housing [LT78a]. hypotheses [Nef79, SG76]. Hypothesis [NS79, Att77, MV74, Spe75, Sta74, Tsu76].

identifiability [Dei78, Gab78]. Identification [AS74, Ger76, Nel74, Koh78, Mon78, Phi73]. II [SR78]. impact [McC74, NN78, UM78]. Implications [Aok74, JP78]. implicit [Gal77, GJ79]. imports [MP78]. improved [Sri73, Ull74]. including [Zel78]. income [Att77, Cha77, Had76, HS76, KvD78, Sta74]. incomes [Afr74]. Incomplete [Dag76, Dag73]. Inconsistency [DG78]. inconsistent [Hen79]. incorporating [MP78]. independent [Aig73d, Ros77]. Index [Ano73b, Ano74c, Ano76c, Ano77e, Ano77f, Ano78e, Ano78f, Ano79k, Ano79l, Ano79m, Die76, Ho79]. Indonesian [Kak77]. industrial [DN73, PW79]. industry [RS77, Wil79]. inefficiency [SL79]. inequality [And74, And77, Gas75, TZ79]. inference [GJ79]. influence [BL75]. Information [Den76, Ame75, Bes79, BW73, DE74, GQ75, Kel75, Rad78, Ric75a, Sch76, Maz75]. inputs [Fus77]. instantaneous [PH79, Pri79]. instrumental [GJ79, Hen79]. instruments [Kle79a, McC74]. insurance [Mor79, NMM^{+}79]. integrated [DM78]. integration [Die75]. interacting [Aok74]. interdependence [GR79b]. International [Gav76, Ric75b, Wei75]. interpretation [FS74]. interval [Bet77]. intervals [Sal76]. introduction [Aig79b, AM79, Ano79h, Hur73, Dag77]. inventory [Tri73]. inverse [Sch73]. invertible [PS77]. investment [Aig73b, CMP77, Sas78, Tri73, Lev77a]. issues [NMM^{+}79]. item [Ano74d, Ano77g, Ano79n]. Items [Ano76d, Ano75e, Ano75f, Ano77h, Ano78g].

J [Lev77a, Mad76, Maz75, O’B75, O’H77, Ric75b]. James [Jud75, Lap76].

Jan [Gav76]. January [Ano77i, Ano79q]. Japanese [Sas78, Tsu76]. Jati [Tak73]. Jenkins [And75]. Johansen [Die75]. joint [GM79a, Nel76, Vin76]. Jorgenson [Mit77]. Journal [Ano79a]. Judge [Aig75]. July [Ano74f, Ano77j]. June [Ano73d, Ano78k, Ano79r].

Karl [Aig73b]. Klein [Has76]. Koerts [Neu77]. Konijn [Tag75]. Kuh [Hur73].

L [Die75, Lap76]. labor [BC73, Cra79]. Labour [Wal78]. lag [Cor77, Gut76, Kas76, SG76, Wei78, ZW73]. lagged [Spe75]. lags [Pag78]. Langaskens [Dag77]. Lansing [Wei75]. large [Mar75]. large-concentration-parameter [Mar75]. law [HR74, HR77]. Lawrence [Wei75]. Least [Sti76, Aig74, Ame74, Ame75, Bes79, Dag73, DS78b, Gal77, HH74, Lea75, Mar76, Phi77a, Saw78, Sch75b]. least-squares [Ame74, Ame75, Bes79, Gal77, Lea75, Phi77a]. l’econométrie [Dag77]. Lectures [Aig73a]. lemma [And75, And77]. Leonard [Mad76]. level [FG78, HV77, HKM79]. levels [TW75]. Librairie [Dag77]. life [Sta74]. light [(Mu77]. likelihood [Ame75, BM78, Ber79, Bes79, Cor79, GR79a, Mag78, Rad78, Sch76]. limited [Ame75, Kel75, LT78a, Rad78]. limited-information [Ame75, Kel75, Rad78]. Linear [Hsi79, Afr74, Cor79, DTH78, De 78, Dei78, DS78a, FHM75, FB74, GJ79, HP74, Hol77, LT75, McE77b, Mon78, Rad78, Sas78, Sch75a, Sch75b, TTC77, Ull74]. linkage [Ric75b]. load [Aig79a, GERA79, Koe79, MMA79, Uri79]. Local [Koh78]. location [WG78]. logit [Ame78, GS79]. London [Lee73, Poi75a]. long [Die75]. loss [Lov73, YJ76, Zel78].

M [Lee73]. macro [Hat78, Pag73a, TT76]. macro-economic [Hat78]. macroeconomic [Aok74, Tri75]. macroeconomics [Hur73]. male [Kie78]. management [Lap76]. managerial [Bal75]. manufacturing [BC73, DM77, Fus77, HV77, UM78]. many [Fus77]. March [Ano73c, Ano77k]. marco [Die75]. Market [Hun79, AR75, Cra79]. markets [GR79b, Gro75]. Markov [GQ73]. Masanao [Cho77]. mathematical [Aig73a, Bal75, Aig73b]. matrix [Bal73, Mag78, Ric75a, Sch76]. Maximum [Mag78, Man75, Ame75, BM78, Ber79, Bes79, Cor79, GR79a, Rad78, Sch76]. maximum-likelihood [Ame75, Bes79, Rad78, Sch76]. May [Ano74e, Ano75h, Ano76f, Ano77n]. Mayer [Sta74]. Mean [Hol77, GR79a, GR74, NS78, Saw73]. means [Zel78]. measurability [Sch73]. Measurement [Afr74, Lev77b, NMM^{+}79, Ger76, SJ75]. measures [Gas75]. meat [CM77]. meeting [Ano78g]. MELO [Zel78]. metals [Wil79]. method [Kak78, Sri73, TZ79]. methodology [Aig79b, HH74, Ken75]. Methods [Tak73, Gil75, Hat78, Lee73, O’B75, Pal77]. Michigan [Wei75]. micro [Die75, TT76]. migration [AGS79]. minimum [UM78, Zel78]. missing [SM75]. misspecification [HC77]. mix [Aok74]. mixed [Gil77, Uri79]. mode [KK74, WG78]. model [Aig73c, AGS79, Ame77a, Ame77b, Ame78, Att77, Ber79, Bus79, But76, CW79, DTH78, De 78, DG78, DS78b, FHM75, Fit73, Gew78, GQ73, GQ75, GDP79, GS79, Gut76, HP74, Hat74, Hat76, HKM79, HF73, Kad76, Kni77, LT75, Mag78, Man75, MS78, Men77, Mor79, NS79, Rei79, RS77, Saw78, Sch75b, Tau79, Wei78, ZW73]. Modeling [Ave79, Nef79, Gro75, PP79]. modelling [Tay79]. models [Aig73b, ALS77, BM78, Bes79, Bet77, BL75, Bre73, Cor79, Dag76, DN73, Dei78, DS78a, DE74, FB74, Ger76, Gus78, Hat78, Hau77, Ho79, Hol77, Ken75, Koh78, KK74, Laf77, LM79, Lan79, LT78a, McE77b, Mon78, MV74, Nef79, Nel77, Pag73b, Pal77, Phi73, Pie75, Plo79, Rob77, SG76, SR75, Szr78, Wes74, WG78, ZP74, Ric75b, Sch75a]. modified [Ame75, Ame77a]. moments [Kni77, MS78, Saw78]. monetary [Ave79, Fro74, McC74, Pag73a]. money [Spi77]. Monograph [Maz75]. monographs [Poi75a]. Monte [DM78, Gil75, HH74, Sow73, YJ76]. monthly [But76]. moving [ASW77, DM78, Fit73, Men77, Nel74, Nel76, NS79, PS77]. MSE [Aig74, McE77b]. Multi [Dan79, KK74]. Multi-period [Dan79]. Multicollinearity [Sch75a]. Multidimensional [Mai78]. multiple [Dag73, Has76, PZ78]. multiple-output [Has76]. multiplicative [God78]. multipliers [GB78]. multivariate [Ame78, FHM75, GS79, Jud75, Ric75a, Szr78].

N [Wil74]. N.Y [Gav76]. national [Ric75b]. natural [GR79a]. natural-conjugate [GR79a]. near [GR77]. Neelman [Sch75a]. neoclassical [App78]. Netherlands [Sch75a]. News [Ano74d, Ano75e, Ano75f, Ano76d, Ano77h, Ano77g, Ano79n]. no [Aig75, Maz75, Poi75a, Sch75a]. Non [Aok74, GJ79, HV77, McE77b, PS77, Rad78]. non-Bayesian [HV77]. Non-interacting [Aok74]. non-invertible [PS77]. non-linear [GJ79, Rad78]. non-spherical [McE77b]. Nonlinear [Laf77, Ame74, Ame75, Bes79, Gal75, Gal77, Hat78, Lee73]. normal [Bal73, Ric75a]. Normalization [Fis76, AS74, Kad78a]. North [Aig73b, Aig75, Bal75, Blo73, Cho74, Cho77, Die75, Hur73, Lee73, Mad76, Mit77, O’H77, Pag73a, Ric75b, Tag75, Tak73, Wil74]. North-Holland [Aig73b, Aig75, Bal75, Blo73, Cho74, Cho77, Die75, Hur73, Lee73, Mad76, Mit77, O’H77, Pag73a, Ric75b, Tag75, Tak73, Wil74]. note [AS74, Ame77b, BL75, BW73, Cor79, Dhr74, Gut76, Lan79, Lov73, Mar76, Rad78, Ric75a, Sch78, Sch73, Sch75b]. November [Ano75j, Ano76h, Ano77m]. null [Spe75]. numbers [Die76, Sch77]. numerical [Saw73].

O [Kad76]. observation [Aig73d, Aig74]. observations [Att77, Dag73, Dag76, Kak77, Sch77, SM75, Poi75a]. obtaining [Vis78]. occupational [Kie78]. October [Ano73e, Ano78j, Ano79s]. Officer [Wei75]. oligopolistic [GR79b]. OLS [DG78, Lev77b]. Omitted [GM79b]. omitting [Vis78]. one [Spe75]. operators [GB78]. Optimal [Kle79a, Koe79, Pap78, Aig79a, Aig79b, FG78, TW75, Cho74, Cho77]. Optimality [DS78b]. optimizing [CW79]. Option [CMP77, Kas76]. options [Pan76]. order [BM78, GK78, Gui74, Men77, Mon78, Nel74, Phi79b]. ordinal [Gre74]. ordinary [HH74]. output [Has76]. overdifferencing [PS77].

P [Aig73b, Lev77a, Tri75, Wil74]. Pages [Ano73f, Ano73d, Ano73c, Ano73e, Ano74h, Ano74f, Ano74e, Ano74g, Ano75i, Ano75g, Ano75h, Ano75j, Ano76g, Ano76e, Ano76f, Ano76h, Ano77i, Ano77j, Ano77k, Ano77n, Ano77m, Ano77l, Ano78l, Ano78h, Ano78m, Ano78i, Ano78k, Ano78j, Ano79o, Ano79u, Ano79t, Ano79q, Ano79r, Ano79s, Ano79p]. parameter [Mar75, MNS78, Phi73, Tsu77, Wei78]. parameters [GM79b, KvD78, Mag78]. Pareto [HR74, HR77]. Parking [WG78]. part [TW79]. partial [DG78]. partially [Kni77, NS78]. patterns [HKP77]. peak [Koe79, MMA79, SB79, Uri79]. peak-load [MMA79]. Penn [ZW73]. period [Dan79]. permanent [Att77, Sta74, Sta74]. petroleum [RS77]. Pierce [Lap76]. Pindyck [Cho74]. pitfalls [Bur75]. Plains [Gav76]. planning [Aig75, Bet77, Cho74]. point [Fis76, Kad78a]. points [GK78]. policy [Aok74, Ave79, Fro74, McC74, Cho74, Mit77, O’H77]. poly [Drè77]. poly-t [Drè77]. Polynomial [GB78, HH78, MP78, Pag78]. population [Zel78]. portfolio [Lap76]. Posterior [PZ78]. Powell [Pag73a]. power [Afr74, HP74, LT75]. pp [Aig73a, Aig73b, Aig75, Bal75, Blo73, Dag77, Hur73, Lee73, Maz75, Mit77, O’B75, Pag73a, Poi75a, Ric75b, Tak73, Tri75, Wil74]. pp. [Cho74, Cho77, Die75, Jud75, Sch75a, Wei75]. pp.xv [Tag75]. Pre [TW79, TW75]. pre-test [TW75]. Pre-testing [TW79]. precision [Aig79c]. Prediction [Lan79, Tau79, Nel74, Sal76]. Predictions [Wes74, GR74, Sal76]. predictors [Pan76]. preferences [CM77, Koe79]. preliminary [BJY73, JYB73, TW75]. presence [SJ75]. Press [Gav76, Jud75, Kad76, Lap76, O’B75, Sta74, Tri75, Maz75, Sch75a]. price [App79, BKK79, Kas76, MP78, Nef79, Pan76]. pricing [Aig79c, Atk79, Dan79, GERA79, HKM79, Koe79, LB79, MMA79]. primary [Wil79]. principal [Had76]. priori [Maz75]. probability [Aig73a, GM79a]. probit [AGS79]. problem [Pap78, Phi73]. problems [Sas78, Wil74]. process [Nel74, PS77]. processes [ASW77, DM78, Nel76, Ros77]. product [Tsu76]. production [ALS77, App78, BL75, CM79, Fus77, HV77, Has76, JP78, LT78b, SL79, TT76, Tsu76, Vin76, Die75]. program [Ano79j]. programming [Tak73]. projecting [DN73]. Properties [JYB73, DTH78, DE74, GS79, Hyl77]. pseudo [Sch73]. pseudo-inverse [Sch73]. Publ [Aig73a, Aig73b, Blo73, Hur73, Lee73, Pag73a, Tak73]. public [AR75, GV77]. Publishing [Mad76]. purchasing [Afr74].

quadratic [Bal73]. qualitative [Ame77a]. qualities [DLA78]. Quandt [Lee73]. quantitative [Ano79j, O’H77, O’B75]. quarterly [But76, Gil77].

R [Cho74, Lee73, Lev77a, Pag73a, Ric75b, Wil74]. railroad [Has76]. random [AGS79, GR77, GDP79, Phi79a]. ratio [Szr78]. Rational [Gro75, Pag78, Hun79]. real [DM77]. reconstruction [Ho79]. Recursions [Phi77a]. reduced [Kni77, NS78]. reduced-form [Kni77]. Regression [Aig73d, Gil77, Bal73, BJY73, BW73, Cor79, Dag73, Drè77, DS78b, GR79a, GR74, GN76, HC77, Hsi79, Hyl77, JYB73, LT75, McE77b, NN78, Nel77, Neu77, Phi77b, Sch78, Sch75b, Sri73, SD79, SM79, Szr78, TTC77, TZ79, Ull74, Vis78, Zel78, Poi75a]. regressions [Bus79, Gal75, GQ73, GN74, McE77a, Sch77, Spe79, SM75]. regressors [Den78, Pie75, PZ78, Spe75]. regulation [Tri75]. Rejoinder [Kad78b]. related [Sta74]. relations [Pag73a]. relationships [Wei75]. Relative [SR75, SR78, DLA78, McC74, SL79]. remainders [SG76]. repeated [GR74]. Reply [Fis78]. reported [HR74]. Reserve [ZW73]. Residential [GERA79, HKP79, LB79, Tay79]. residuals [DS78a]. response [Ame77a, CW79, MP78]. Responsiveness [Atk79]. restricted [Kni77, Lea75, NS78]. restrictions [BPR77, Dei78, Den78, Hol77, JYB73, McE77b]. result [Lea75]. results [Atk79, Men77]. Retail [Tri73]. Retention [Gun74]. returns [CMP77, GV77]. rev [O’H77]. Review [Aig73a, Aig73b, Aig75, Bal75, Blo73, Cho74, Cho77, Dag77, Die75, Gav76, Hur73, Jud75, Kad76, Lap76, Lee73, Lev77a, Mad76, Maz75, Mit77, O’B75, O’H77, Pag73a, Poi75a, Ric75b, Sch75a, Sta74, Tag75, Tak73, Tri75, Wei75, Wil74]. revisited [And77, Cha77, Has76]. Richard [Blo73, Hur73]. ridge [Vin76]. right [Vis78]. Rinehart [Jud75]. Rodan [Wil74]. Rosentein [Wil74]. Rosentein-Rodan [Wil74]. Rothenberg [Maz75]. Rotterdam [O’B75, Tri75]. round [Ame77a]. Rowley [Lev77a]. rule [Aig79a, YJ76]. run [Aig73c, Die75].

S [Cho74, Jud75, Kad76, Lee73, Mad76, Tag75, Wil74]. same [Den78]. Sample [Aig79c, Aig79a, GS79, HH74, Hyl77, Phi77b, Spe75, Tag75]. samples [Hur79]. sampling [Bre73, Phi79b, Ull74]. Savage [Mad76]. scaling [Mai78]. schemes [MNS78]. Schmalensee [Blo73, Hur73]. school [AR75]. Sciences [Gav76, Bal75, Kad76]. score [Man75]. seasonal [Plo79]. seasonally [PP79]. second [Ame77a, BM78, NMM^{+}79]. second-order [BM78]. second-round [Ame77a]. section [BKK79]. sector [DM77, HV77]. Seemingly [Gal75, Bus79, DS78b, McE77a, Phi77b, Sch77, Sch78, Spe79, Sri73, SD79, SM75]. selection [Mor79]. Sellekaerts [Gav76]. Seminar [Kad76]. Sengupta [O’H77, Tak73]. separability [BPR77, Woo78]. September [Ano74g, Ano77l, Ano79p]. serial [NN78, Spe75, Szr78]. serially [SR78]. series [And74, DLA78, FHM75, Had76, Koh78, NG79, NS79, Pal77, Pap78, PP79, ZP74]. series/econometric [Uri79]. sets [Den78]. Several [Hat76]. share [Woo79]. Shell [Aig73b]. shift [FHM75, Tsu77]. short [Aig73c, Die75]. short-run [Aig73c]. side [Nef79]. sign [Lea75, Vis78]. significance [BJY73, FG78, GK78, JYB73, NN78, TW75]. simple [MS78, SR75, SR78]. simultaneous [Bes79, CN77, Con79, Dag76, Den76, DE74, Gal77, GJ79, Ger76, Gew78, Hat76, Hau77, Hen76, Kel75, Kni77, MS78, MV74, Pal77, Rei79, SW78, Spi77, ZP74]. simultaneous-equation [Kni77]. simultaneous-equations [Bes79, Dag76]. Single [Den78]. Single-equation [Den78]. singular [Hol77]. size [Aig79a]. slopes [FHM75]. small [GS79, HH74, Kle79a, Spe75]. Smooth [Cor77]. smoothing [Cor77]. Social [FH79, Bal75, Kad76, NMM^{+}79]. Society [Ano78g]. Solow [Gut76]. Some [BJY73, Bre73, CM79, FHM75, GS79, Mai78, Mar75, McC74, Men77, TTC77, HP74, LT78a, MS78, SM75, SR75, SR78]. space [Aig75]. spaces [Cor77]. special [Sch76]. Specification [MP78, Bet77, Bur75, (Mu77, Gew78, Gui74, JP78, Woo79]. Spectral [GV77]. spectrum [Hyl77]. spherical [McE77b]. spline [Cor77]. splines [Poi75b]. Spurious [GN74]. square [GR74, Hol77, Saw73]. squared [GR79a, NS78, YJ76]. squares [Aig74, Ame74, Ame75, Bes79, Dag73, DS78b, Gal77, HH74, Lea75, Mar76, Phi77a, Saw78, Sch75b, Sti76]. stability [Gus78]. stabilization [Cho74]. stage [Ame74, Ame75, Bes79, Gal77, HH74, Mar76, Phi77a]. state [Wei75]. stationary [PP79]. statistic [DTH78]. Statistical [GJ79, FS74, Poi75a, Bal75, Tag75]. Statistics [Mad76, GS79, Aig73a]. steel [Tsu76]. Stein [YJ76]. Stein-rule [YJ76]. step [Ame78, Dhr74, Hat74, Hat76]. Stephan [Aig73a]. stipends [CMP77]. Stochastic [JP78, Woo79, ALS77, Dan79, JYB73, LT78b, Man75, Nel77, Pie75, SL79, Sti76, Tak73]. stock [Pan76]. strong [Koh78]. structural [Dei78, Zel78, Kad76]. structure [ASW77, FB74, Hen76, Kas76, McC74]. structure-based [McC74]. structures [BC73, Fus77]. Studies [Wil74, Bal75, Mit77, Pag73a, Sch75a, Sow73, Wei75, Aig75, Mad76]. study [AGS79, DM78, Gun74, Gus78, Has76, Hyl77, Kle79b, Mor79, NMM^{+}79, Wal78, WG78]. subject [Aig73d, Gro75, TTC77]. subsidized [Kie78]. substitution [BC73, SJ75]. superlative [Die76]. supply [Spi77, Wal78, Wei75]. surfaces [CW79]. survey [PH77, SD79, Tag75]. switching [GQ73, TZ79]. system [Cho77, Gal77, GJ79, Spe79, Spi77, Uri79]. systematic [Bre73]. systems [CN77, Hen79, Kel75, Kle79b, MP78, Nel76, PH77, Sch78]. Szego [Aig73b].

T [Sta74, Drè77]. Takashi [Aig75]. Takayama [Aig75]. taking [App79]. teachers [AR75]. Technical [Wil79, SL79]. technique [HH78]. technologies [Bur75]. temporal [BKK79, Bre73, PH77, Wei78]. temporally [Hsi79]. terms [Pag73b]. test [FG78, Fro74, GK78, GS79, HKM79, NN78, TW75, Tsu76, Tsu77]. Testing [App78, App79, Fit73, Gew78, God78, Gus78, Had76, HC77, BPR77, MV74, NS79, SW78, TW79, Woo78]. tests [BJY73, DLA78, FHM75, Gui74, HP74, Hol77, JYB73, LT75, McE77b, NN78, SG76, Spe75, Szr78]. Theil [Bal75]. Theoretical [PW79]. theories [Sta74]. Theory [Hun79, Aig73a, App78, Bur75, Cho74, Cho77, Gav76, O’H77, Sta74, Tag75]. there [Hur79]. Thomas [Maz75]. Thorbecke [O’H77]. Three [Gal77, Bes79, Mar76]. Three-stage [Gal77, Bes79, Mar76]. thresholds [Nel77]. Tilburg [Sch75a]. Time [ZP74, Aig79c, And74, Atk79, DLA78, FHM75, GERA79, HKM79, Koe79, Koh78, KK74, LB79, NG79, NS79, Pal77, PW79, Pap78, PP79, Phi73, Rob77, Tay79, Uri79, Wal78, Aig75]. time-additive [Koe79]. time-of-day [Aig79c, Atk79, GERA79, HKM79, LB79, Tay79]. time-series [Uri79]. time-series/econometric [Uri79]. Tinbergen [Gav76]. traditional [YJ76]. trainees [Gun74, Kie78]. training [Kie78]. transformation [GN76, NG79]. transit [WG78]. translog [BC73, CM79]. travel [KK74]. treatments [SG76]. Trivedi [Lev77a]. true [Spe75]. truncated [Hur79]. truncation [SG76]. TSLS [Saw73]. two [Ame74, Ame75, Ame78, Dhr74, HV77, Hat74, Hat76, HKM79, HH74, Phi77a, SM79]. two-level [HV77, HKM79]. two-stage [Ame74, Ame75, HH74, Phi77a]. two-step [Ame78, Dhr74, Hat74, Hat76]. type [DTH78]. typical [DLA78].

U.S [Bal75, Mit77, O’B75]. U.S. [BC73, CM77, MP78, Wil79]. unbiased [Bal73]. uncertainty [BL75, Gro75]. unconditional [GR74]. Uncorrelated [DS78a]. under-reported [HR74]. underlying [Nef79]. underreported [HR77]. unequal [Sch77]. univariate [Pal77]. University [Lap76, O’B75, Sta74, Tri75, Maz75, Sch75a, Wei75]. unknown [Mag78]. unobservable [Att77]. unobserved [Ave79, Nel77]. unrelated [Bus79, DS78b, Gal75, McE77a, Phi77b, Sch77, Sch78, Spe79, Sri73, SD79, SM75]. unstable [Gus78]. upon [NN78]. use [But76, Dag73, GN76, Poi75b, Sal76]. used [JYB73]. using [Att77, Drè77, Gil77, Hsi79, KK74, NG79, Spi77, TZ79]. utility [CM77, GV77, Man75].

value [DM77, GQ75, KK74]. value-added [DM77]. values [CMP77]. variability [GM79b]. variable [Aig73d, Ave79, GJ79, LT78a, Spe75, Vis78]. variables [Aok74, Att77, GN76, Gun74, GM79b, Hat78, Hau77, Hen79, Sal76]. variance [Bal73, Laf77, Szr78, Tau79, TW75]. variance-components [Tau79]. variance-covariance [Bal73]. variance-ratio [Szr78]. variation [MNS78]. various [SG76]. vector [Gui74, Phi79a]. viii [Maz75, O’B75, Pag73a, Sch75a]. Vishwakarma [Tri75]. vol [Bal75, Blo73].

W [Gav76, Mit77]. wage [Ken75, UM78]. walks [GR77]. Wallis [GK78]. Watson [FG78]. weak [Woo78]. Weaker [McE77b]. wealth [Sta74]. White [Gav76]. Wiley [Lev77a]. Williams [Pag73a]. Winston [Jud75].

x [Poi75a, Wei75]. xi [Aig73a, Tak73, Tri75]. xii [Aig75, Cho74, Lee73, Ric75b, Wil74]. xiii [Blo73]. xiv [Cho77]. xix [Jud75]. xvi [Bal75].

Yale [Lap76, Maz75]. York [Aig73a, Jud75, Kad76, Lev77a, Maz75]. Yvan [Dag77].

Zellner [Mad76, Phi77b]. Zubrzycki [Aig73a].

References

A:1973:E

[ADZ73] D. J. A., P. J. D., and A. Z. Editorial. Journal of Econometrics, 1(1):1, March 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407673900018.

Afriat:1974:MPP

[Afr74] S. N. Afriat. Measurement of the purchasing power of incomes with linear expansion data. Journal of Econometrics, 2(4):343– 364, December 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900190.

Akin:1979:RCP

[AGS79] John S. Akin, David K. Guilkey, and Robin Sickles. A random coefficient probit model with an application to a study of migration. Journal of Econometrics, 11(2–3):233–246, October/ December 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900381.

Aigner:1973:BRL

[Aig73a] D. J. Aigner. Book review: Lectures in probability theory and mathematical statistics: Stephan Zubrzycki, (Elsevier Publ. Co., Amsterdam and New York, 1972) xi + 321 pp. (Dfl. 47.50). Journal of Econometrics, 1(4):397, December 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407673900249.

Aigner:1973:BRM

[Aig73b] D. J. Aigner. Book review: Mathematical models in investment and finance: G. P. Szego and Karl Shell (eds)., (North-Holland Publ. Co., Amsterdam, 1973) 656 pp. ($41.50). Journal of Econometrics, 1(4):397–398, December 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407673900250.

Aigner:1973:EEM

[Aig73c] Dennis J. Aigner. On estimation of an econometric model of short-run bank behaviour. Journal of Econometrics, 1(3):201– 228, October 1973. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407673900080.

Aigner:1973:RBI

[Aig73d] Dennis J. Aigner. Regression with a binary independent variable subject to errors of observation. Journal of Econometrics, 1(1): 49–59, March 1973. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407673900055.

Aigner:1974:MDL

[Aig74] D. J. Aigner. MSE dominance of least squares with errors-of-observation. Journal of Econometrics, 2(4):365–372, December 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407674900207.

Aigner:1975:BRS

[Aig75] Dennis J. Aigner. Book review: Studies in economic planning over space and time, contributions to economic analysis no. 82: George Judge and Takashi Takayama, (North-Holland, Amsterdam, 1973) xii + 727 pp., $72.00. Journal of Econometrics, 3(1): 94, February 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900706.

Aigner:1979:BAO

[Aig79a] Dennis J. Aigner. Bayesian analysis of optimal sample size and a best decision rule for experiments in direct load control. Journal of Econometrics, 9(1–2):209–221, January 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679901040.

Aigner:1979:BIM

[Aig79b] Dennis J. Aigner. A brief introduction to the methodology of optimal experimental design. Journal of Econometrics, 11(1):7– 26, September 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900514.

Aigner:1979:SDE

[Aig79c] Dennis J. Aigner. Sample design for electricity pricing experiments: Anticipated precision for a time-of-day pricing experiment. Journal of Econometrics, 11(1):195–205, September 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900587.

Aigner:1977:FES

[ALS77] Dennis Aigner, C. A. Knox Lovell, and Peter Schmidt. Formulation and estimation of stochastic frontier production function models. Journal of Econometrics, 6(1):21–37, July 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900525.

Aigner:1979:EI

[AM79] Dennis J. Aigner and Carl N. Morris. Editors’ introduction. Journal of Econometrics, 11(1):1–5, September 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679900502.

Amemiya:1974:NTS

[Ame74] Takeshi Amemiya. The nonlinear two-stage least-squares estimator. Journal of Econometrics, 2(2):105–110, July 1974. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407674900335.

Amemiya:1975:NLI

[Ame75] Takeshi Amemiya. The nonlinear limited-information maximumlikelihood estimator and the modified nonlinear two-stage leastsquares estimator. Journal of Econometrics, 3(4):375–386, November 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407675900548.

Amemiya:1977:MSR

[Ame77a] Takeshi Amemiya. The modified second-round estimator in the general qualitative response model. Journal of Econometrics, 5(3):

295–299, May 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900409.

Amemiya:1977:NHM

[Ame77b] Takeshi Amemiya. A note on a heteroscedastic model. Journal of Econometrics, 6(3):365–370, November 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900069.

Amemiya:1978:TSE

[Ame78] Takeshi Amemiya. On a two-step estimation of a multivariate logit model. Journal of Econometrics, 8(1):13–21, August 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407678900866.

Anderson:1974:ITS

[And74] O. D. Anderson. An inequality with a time series application. Journal of Econometrics, 2(2):189–193, July 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407674900396.

Anderson:1975:LAB

[And75] O. D. Anderson. On a lemma associated with Box, Jenkins and granger. Journal of Econometrics, 3(2):151–156, May 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407675900433.

Anderson:1977:ILR

[And77] Oliver D. Anderson. An inequality and a lemma revisited. Journal of Econometrics, 6(1):135–140, July 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900598.

Anonymous:1973:A

[Ano73a] Anonymous. Acknowledgement. Journal of Econometrics, 1(4): 407, December 1973. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407673900316.

Anonymous:1973:I

[Ano73b] Anonymous. Index. Journal of Econometrics, 1(4):409–410, December 1973. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407673900328.

Anonymous:1973:PM

[Ano73c] Anonymous. Pages 1–113 (March 1973). Journal of Econometrics, 1(1):??, March 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1973:PJ

[Ano73d] Anonymous. Pages 115–200 (June 1973). Journal of Econometrics, 1(2):??, June 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1973:PO

[Ano73e] Anonymous. Pages 201–316 (October 1973). Journal of Econometrics, 1(3):??, October 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1973:PD

[Ano73f] Anonymous. Pages 317–410 (December 1973). Journal of Econometrics, 1(4):??, December 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1974:A

[Ano74a] Anonymous. Acknowledgement. Journal of Econometrics, 2(4): 395, December 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900244.

Anonymous:1974:EB

[Ano74b] Anonymous. Editorial Board. Journal of Econometrics, 2(1): ifc, May 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900268.

Anonymous:1974:I

[Ano74c] Anonymous. Index. Journal of Econometrics, 2(4):397–398, December 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900256.

Anonymous:1974:NI

[Ano74d] Anonymous. News item. Journal of Econometrics, 2(3):305–306, September 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900086.

Anonymous:1974:PM

[Ano74e] Anonymous. Pages 1–104 (May 1974). Journal of Econometrics, 2(1):??, May 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1974:PJ

[Ano74f] Anonymous. Pages 105–198 (July 1974). Journal of Econometrics, 2(2):??, July 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1974:PS

[Ano74g] Anonymous. Pages 199–306 (September 1974). Journal of Econometrics, 2(3):??, September 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1974:PD

[Ano74h] Anonymous. Pages 307–398 (December 1974). Journal of Econometrics, 2(4):??, December 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1975:Aa

[Ano75a] Anonymous. Acknowledgement. Journal of Econometrics, 3(4): 407–408, November 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407675900597.

Anonymous:1975:Ab

[Ano75b] Anonymous. Articles. Journal of Econometrics, 3(4):409–410, November 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900603.

Anonymous:1975:EB

[Ano75c] Anonymous. Editorial Board. Journal of Econometrics, 3(1): ifc, February 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900615.

Anonymous:1975:E

[Ano75d] Anonymous. Erratum. Journal of Econometrics, 3(4):405, November 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900585.

Anonymous:1975:NIa

[Ano75e] Anonymous. News items. Journal of Econometrics, 3(1):103– 104, February 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900767.

Anonymous:1975:NIb

[Ano75f] Anonymous. News items. Journal of Econometrics, 3(3):321– 323, August 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900408.

Anonymous:1975:PF

[Ano75g] Anonymous. Pages 1–104 (February 1975). Journal of Econometrics, 3(1):??, February 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1975:PM

[Ano75h] Anonymous. Pages 105–203 (May 1975). Journal of Econometrics, 3(2):??, May 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1975:PA

[Ano75i] Anonymous. Pages 205–323 (August 1975). Journal of Econometrics, 3(3):??, August 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1975:PN

[Ano75j] Anonymous. Pages 325–410 (November 1975). Journal of Econometrics, 3(4):??, November 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1976:A

[Ano76a] Anonymous. Acknowledgement. Journal of Econometrics, 4(4): 401–402, November 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407676900300.

Anonymous:1976:EB

[Ano76b] Anonymous. Editorial board. Journal of Econometrics, 4(1): ifc, February 1976. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407676900130.

Anonymous:1976:I

[Ano76c] Anonymous. Index. Journal of Econometrics, 4(4):403, November 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407676900312.

Anonymous:1976:NI

[Ano76d] Anonymous. News items. Journal of Econometrics, 4(3):301, August 1976. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407676900403.

Anonymous:1976:PF

[Ano76e] Anonymous. Pages 1–99 (February 1976). Journal of Econometrics, 4(1):??, February 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1976:PM

[Ano76f] Anonymous. Pages 101–204 (May 1976). Journal of Econometrics, 4(2):??, May 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1976:PA

[Ano76g] Anonymous. Pages 205–301 (August 1976). Journal of Econometrics, 4(3):??, August 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1976:PN

[Ano76h] Anonymous. Pages 303–403 (November 1976). Journal of Econometrics, 4(4):??, November 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1977:A

[Ano77a] Anonymous. Acknowledgement. Journal of Econometrics, 6(3): 395–396, November 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407677900100.

Anonymous:1977:EBa

[Ano77b] Anonymous. Editorial Board. Journal of Econometrics, 5(1): iv, January 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900306.

Anonymous:1977:EBb

[Ano77c] Anonymous. Editorial Board. Journal of Econometrics, 6(1): ifc, July 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900501.

Anonymous:1977:E

[Ano77d] Anonymous. Erratum. Journal of Econometrics, 6(2):261, September 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900197.

Anonymous:1977:Ia

[Ano77e] Anonymous. Index. Journal of Econometrics, 5(3):405–406, May 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407677900495.

Anonymous:1977:Ib

[Ano77f] Anonymous. Index. Journal of Econometrics, 6(3):397–398, November 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900112.

Anonymous:1977:NIb

[Ano77g] Anonymous. News item. Journal of Econometrics, 5(3):403, May 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407677900483.

Anonymous:1977:NIa

[Ano77h] Anonymous. News items. Journal of Econometrics, 5(2):261– 263, March 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440767790029X.

Anonymous:1977:PJa

[Ano77i] Anonymous. Pages 1–133 (January 1977). Journal of Econometrics, 5(1):??, January 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1977:PJb

[Ano77j] Anonymous. Pages 1–145 (July 1977). Journal of Econometrics, 6(1):??, July 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1977:PMa

[Ano77k] Anonymous. Pages 135–263 (March 1977). Journal of Econometrics, 5(2):??, March 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1977:PS

[Ano77l] Anonymous. Pages 147–261 (September 1977). Journal of Econometrics, 6(2):??, September 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1977:PN

[Ano77m] Anonymous. Pages 263–398 (November 1977). Journal of Econometrics, 6(3):??, November 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1977:PMb

[Ano77n] Anonymous. Pages 265–406 (May 1977). Journal of Econometrics, 5(3):??, May 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1978:A

[Ano78a] Anonymous. Announcement. Journal of Econometrics, 7(3): 397, April 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900635.

Anonymous:1978:C

[Ano78b] Anonymous. Corrigenda. Journal of Econometrics, 8(2):265, October 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900362.

Anonymous:1978:EBa

[Ano78c] Anonymous. Editorial board. Journal of Econometrics, 7(1): ii, February 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900015.

Anonymous:1978:EBb

[Ano78d] Anonymous. Editorial Board. Journal of Econometrics, 8(1): ifc, August 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900842.

Anonymous:1978:Ia

[Ano78e] Anonymous. Index. Journal of Econometrics, 7(3):399–400, April 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407678900647.

Anonymous:1978:Ib

[Ano78f] Anonymous. Index. Journal of Econometrics, 8(3):399–400, December 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900544.

Anonymous:1978:NIE

[Ano78g] Anonymous. New items: European meeting of the Econometric Society, Geneva, 1978. Journal of Econometrics, 7(1):131– 132, February 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900143.

Anonymous:1978:PAb

[Ano78h] Anonymous. Pages 1–125 (August 1978). Journal of Econometrics, 8(1):??, August 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1978:PF

[Ano78i] Anonymous. Pages 1–132 (February 1978). Journal of Econometrics, 7(1):??, February 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1978:PO

[Ano78j] Anonymous. Pages 127–265 (October 1978). Journal of Econometrics, 8(2):??, October 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1978:PJ

[Ano78k] Anonymous. Pages 133–261 (June 1978). Journal of Econometrics, 7(2):??, June 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1978:PAa

[Ano78l] Anonymous. Pages 263–400 (April 1978). Journal of Econometrics, 7(3):??, April 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1978:PD

[Ano78m] Anonymous. Pages 267–400 (December 1978). Journal of Econometrics, 8(3):??, December 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1979:AJE

[Ano79a] Anonymous. Acknowledgement: Journal of Econometrics — 1978. Journal of Econometrics, 9(1–2):239–240, January 1979. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407679901064.

Anonymous:1979:A

[Ano79b] Anonymous. Announcement. Journal of Econometrics, 10(1): 125, April 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900721.

Anonymous:1979:Ca

[Ano79c] Anonymous. Corrigenda. Journal of Econometrics, 10(2):261, June 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900113.

Anonymous:1979:Cb

[Ano79d] Anonymous. Corrigenda. Journal of Econometrics, 10(3):385, August 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900903.

Anonymous:1979:EBa

[Ano79e] Anonymous. Editorial Board. Journal of Econometrics, 9(1–2): ii, January 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900927.

Anonymous:1979:EBb

[Ano79f] Anonymous. Editorial Board. Journal of Econometrics, 10(1): ii, April 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900599.

Anonymous:1979:EBc

[Ano79g] Anonymous. Editorial Board. Journal of Econometrics, 11(1): ii, September 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900496.

Anonymous:1979:EI

[Ano79h] Anonymous. Editors’ introduction. Journal of Econometrics, 9 (1–2):1–9, January 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407679900939.

Anonymous:1979:E

[Ano79i] Anonymous. Erratum. Journal of Econometrics, 9(3):391, February 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900824.

Anonymous:1979:EDP

[Ano79j] Anonymous. European doctoral program in quantitative economics. Journal of Econometrics, 9(3):393, February 1979. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407679900836.

Anonymous:1979:Ia

[Ano79k] Anonymous. Index. Journal of Econometrics, 9(3):395–396, February 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900848.

Anonymous:1979:Ib

[Ano79l] Anonymous. Index. Journal of Econometrics, 10(3):387–388, August 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900915.

Anonymous:1979:Ic

[Ano79m] Anonymous. Index. Journal of Econometrics, 11(2–3):369, October/December 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900484.

Anonymous:1979:NI

[Ano79n] Anonymous. News item. Journal of Econometrics, 11(2–3):367, October/December 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407679900472.

Anonymous:1979:PAa

[Ano79o] Anonymous. Pages 1–125 (April 1979). Journal of Econometrics, 10(1):??, April 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1979:PS

[Ano79p] Anonymous. Pages 1–205 (September 1979). Journal of Econometrics, 11(1):??, September 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1979:PJa

[Ano79q] Anonymous. Pages 1–240 (January 1979). Journal of Econometrics, 9(1–2):??, January 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1979:PJb

[Ano79r] Anonymous. Pages 127–261 (June 1979). Journal of Econometrics, 10(2):??, June 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1979:POD

[Ano79s] Anonymous. Pages 207–369 (October–December 1979). Journal of Econometrics, 11(2–3):??, October/December 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1979:PF

[Ano79t] Anonymous. Pages 241–396 (February 1979). Journal of Econometrics, 9(3):??, February 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:1979:PAb

[Ano79u] Anonymous. Pages 263–388 (August 1979). Journal of Econometrics, 10(3):??, August 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Aoki:1974:NIC

[Aok74] Masanao Aoki. Non-interacting control of macroeconomic variables: Implications on policy mix considerations. Journal of Econometrics, 2(3):261–281, September 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407674900050.

Appelbaum:1978:TNP

[App78] Elie Appelbaum. Testing neoclassical production theory. Journal of Econometrics, 7(1):87–102, February 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900076.

Appelbaum:1979:TPT

[App79] Elie Appelbaum. Testing price taking behavior. Journal of Econometrics, 9(3):283–294, February 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679900757.

Antos:1975:DMP

[AR75] Joseph R. Antos and Sherwin Rosen. Discrimination in the market for public school teachers. Journal of Econometrics, 3(2): 123–150, May 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900421.

Aigner:1974:INN

[AS74] D. J. Aigner and T. Sawa. Identification and normalization: A note. Journal of Econometrics, 2(4):389–391, December 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407674900220.

Ansley:1977:SMA

[ASW77] Craig F. Ansley, W. Allen Spivey, and William J. Wrobleski. On the structure of moving average processes. Journal of Econometrics, 6(1):121–134, July 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407677900586.

Atkinson:1979:RTD

[Atk79] Scott E. Atkinson. Responsiveness to time-of-day electricity pricing: First empirical results. Journal of Econometrics, 9(1–2):79– 95, January 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900976.

Attfield:1977:EMC

[Att77] Clifford L. F. Attfield. Estimation of a model containing unobservable variables using grouped observations: An application to the permanent income hypothesis. Journal of Econometrics, 6(1): 51–63, July 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900549.

Avery:1979:MMP

[Ave79] Robert B. Avery. Modeling monetary policy as an unobserved variable. Journal of Econometrics, 10(3):291–311, August 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900861.

Aigner:1978:E

[AZ78] Dennis Aigner and Arnold Zellner. Editorial. Journal of Econometrics, 8(3):267, December 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407678900477.

Balestra:1973:BQU

[Bal73] Pietro Balestra. Best quadratic unbiased estimators of the variancecovariance matrix in normal regression. Journal of Econometrics, 1(1):17–28, March 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407673900031.

Balestra:1975:BRS

[Bal75] Pietro Balestra. Book review: Statistical decomposition analysis with applications in the social and administrative sciences: Henri Theil, studies in mathematical and managerial economics, vol. 14 (North-Holland, Amsterdam, 1972) xvi + 337 pp., U.S. $22.50. Journal of Econometrics, 3(3):319, August 1975. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440767590038X.

Berndt:1973:TFS

[BC73] Ernst R. Berndt and Laurits R. Christensen. The translog function and the substitution of equipment, structures, and labor in U.S. manufacturing 1929–68. Journal of Econometrics, 1(1):81– 113, March 1973. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407673900079.

Berzeg:1979:ECM

[Ber79] Korhan Berzeg. The error components model: Conditions for the existence of the maximum likelihood estimates. Journal of Econometrics, 10(1):99–102, April 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679900678.

Besley:1979:CCF

[Bes79] David A. Besley. On the computational competitiveness of full-information maximum-likelihood and three-stage least-squares in the estimation of nonlinear, simultaneous-equations models. Journal of Econometrics, 9(3):315–342, February 1979. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407679900770.

Betancourt:1977:CPI

[Bet77] Roger R. Betancourt. On the consequences of planning interval specification error for the estimation of dynamic models. Journal of Econometrics, 6(2):237–242, September 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900173.

Battese:1975:E

[BF75] George E. Battese and Wayne A. Fuller. Erratum. Journal of Econometrics, 3(2):203, May 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407675900512.

Bock:1973:SCE

[BJY73] M. E. Bock, G. G. Judge, and T. A. Yancey. Some comments on estimation in regression after preliminary tests of significance. Journal of Econometrics, 1(2):191–200, June 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407673900171.

Barth:1979:TCS

[BKK79] James Barth, Arthur Kraft, and John Kraft. A temporal crosssection approach to the price equation. Journal of Econometrics, 11(2–3):335–351, October/December 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679900447.

Blair:1975:NIU

[BL75] Roger D. Blair and Rafael Lusky. A note on the influence of uncertainty on estimation of production function models. Journal of Econometrics, 3(4):391–394, November 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407675900561.

Bloch:1973:BRE

[Blo73] Harry Bloch. Book review: The economics of advertising, Contributions to economic analysis, vol. 80: Richard Schmalensee (North-Holland Publ. Co., Amsterdam, 1972) xiii + 312 pp. ($19.00). Journal of Econometrics, 1(4):401–402, December 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407673900286.

Beach:1978:FML

[BM78] Charles M. Beach and James G. MacKinnon. Full maximum likelihood estimation of second-order autoregressive error models. Journal of Econometrics, 7(2):187–198, June 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900684.

Blackorby:1977:TSR

[BPR77] Charles Blackorby, Daniel Primont, and R. Robert Russell. On testing separability restrictions with flexible functional forms. Journal of Econometrics, 5(2):195–209, March 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900240.

Brewer:1973:SCT

[Bre73] K. R. W. Brewer. Some consequences of temporal aggregation and systematic sampling for ARMA and ARMAX models. Journal of Econometrics, 1(2):133–154, June 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407673900158.

Burgess:1975:DTP

[Bur75] David F. Burgess. Duality theory and pitfalls in the specification of technologies. Journal of Econometrics, 3(2):105– 121, May 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440767590041X.

Buse:1979:GFS

[Bus79] A. Buse. Goodness-of-fit in the seemingly unrelated regressions model: A generalization. Journal of Econometrics, 10(1):109– 113, April 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900691.

Butter:1976:UMQ

[But76] F. A. G. Den Butter. The use of monthly and quarterly data in an ARMA model. Journal of Econometrics, 4(4):311–324, November 1976. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407676900221.

Brook:1973:NEI

[BW73] Richard Brook and T. D. Wallace. A note on extraneous information in regression. Journal of Econometrics, 1(3):315–316, October 1973. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407673900134.

Chamberlain:1977:EIA

[Cha77] Gary Chamberlain. Education, income, and ability revisited. Journal of Econometrics, 5(2):241–257, March 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900276.

Chow:1974:BRO

[Cho74] Gregory C. Chow. Book review: Optimal planning for economic stabilization: The application of control theory to stabilization policy: R. S. Pindyck (North-Holland, Amsterdam, 1973) xii + 167 pp. Journal of Econometrics, 2(2):197–198, July 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407674900414.

Chow:1977:BRO

[Cho77] Gregory C. Chow. Book review: Optimal control and system theory in dynamic economic analysis: Masanao Aoki, (North-Holland, Amsterdam,1976) xiv + 400 pp. Journal of Econometrics, 6(1): 143–144, July 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900628.

Christensen:1977:EUC

[CM77] Laurits R. Christensen and Marilyn E. Manser. Estimating U.S. consumer preferences for meat with a flexible utility function. Journal of Econometrics, 5(1):37–53, January 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900331.

Corbo:1979:TPF

[CM79] Vittorio Corbo and Patricio Meller. The translog production function: Some evidence from establishment data. Journal of Econometrics, 10(2):193–199, June 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407679900046.

Comay:1977:OVS

[CMP77] Yochanan P. Comay, Arie Melnik, and Moshe A. Pollatschek. Option values, stipends and the returns to educational investment. Journal of Econometrics, 6(2):243–260, September 1977. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407677900185.

Carter:1977:CCS

[CN77] Richard A. L. Carter and Anirudh L. Nagar. Coefficients of correlation for simultaneous equation systems. Journal of Econometrics, 6(1):39–50, July 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407677900537.

Conlisk:1979:DSE

[Con79] John Conlisk. Design for simultaneous equations. Journal of Econometrics, 11(1):63–76, September 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440767990054X.

Corradi:1977:SDL

[Cor77] Corrado Corradi. Smooth distributed lag estimators and smoothing spline functions in Hilbert spaces. Journal of Econometrics, 5(2):211–219, March 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407677900252.

Corradi:1979:NCM

[Cor79] Corrado Corradi. A note on the computation of maximum likelihood estimates in linear regression models with autocorrelated errors. Journal of Econometrics, 11(2–3):303–317, October/December 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900423.

Crawford:1979:ELM

[Cra79] Robert G. Crawford. Expectations and labor market adjustments. Journal of Econometrics, 11(2–3):207–232, October/ December 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440767990037X.

Conlisk:1979:MOE

[CW79] John Conlisk and Harold Watts. A model for optimizing experimental designs for estimating response surfaces. Journal of Econometrics, 11(1):27–42, September 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679900526.

Dagenais:1973:UIO

[Dag73] Marcel G. Dagenais. The use of incomplete observations in multiple regression analysis: A generalized least squares approach. Journal of Econometrics, 1(4):317–328, December 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407673900183.

Dagenais:1976:IOS

[Dag76] Marcel G. Dagenais. Incomplete observations and simultaneousequations models. Journal of Econometrics, 4(3):231–241, August 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407676900348.

Dagenais:1977:BRI

[Dag77] Denyse L. Dagenais. Book review: Introduction à l’econométrie: Yvan Langaskens, (Librairie Droz, Geneva, 1975) approx. 670 pp. $30.00. Journal of Econometrics, 6(1):144–145, July 1977. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440767790063X.

Dansby:1979:MPP

[Dan79] Robert E. Dansby. Multi-period pricing with stochastic demand. Journal of Econometrics, 9(1–2):223–237, January 1979. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407679901052.

DeJong:1978:DFF

[De 78] Piet De Jong. Determining the final form of a linear dynamic econometric model. Journal of Econometrics, 8(2):181– 192, October 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900271.

Dhrymes:1974:APF

[DE74] Phoebus J. Dhrymes and H. Erlat. Asymptotic properties of full information estimators in dynamic autoregressive simultaneous equation models. Journal of Econometrics, 2(3):247–259, September 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900049.

Deistler:1978:SIL

[Dei78] Manfred Deistler. The structural identifiability of linear models with autocorrelated errors in the case of cross-equation restrictions. Journal of Econometrics, 8(1):23–31, August 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900878.

Dent:1976:ICS

[Den76] Warren Dent. Information and computation in simultaneous equations estimation. Journal of Econometrics, 4(1):89–95, February 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440767690018X.

Denton:1978:SEE

[Den78] Frank T. Denton. Single-equation estimators and aggregation restrictions when equations have the same sets of regressors. Journal of Econometrics, 8(2):173–179, October 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440767890026X.

Doran:1978:IOE

[DG78] H. E. Doran and W. E. Griffiths. Inconsistency of the OLS estimator of the partial adjustment-adaptive expectations model. Journal of Econometrics, 7(2):133–146, June 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900659.

Dhrymes:1974:NET

[Dhr74] Phoebus J. Dhrymes. A note on an efficient two-step estimator. Journal of Econometrics, 2(3):301–304, September 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407674900074.

Diewert:1975:BRP

[Die75] W. E. Diewert. Book review: Production functions: An integration of micro and marco, short run and long run aspects: L. Johansen, (North-Holland, Amsterdam, 1972) 274 pp. Journal of Econometrics, 3(1):97–99, February 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407675900743.

Diewert:1976:ESI

[Die76] W. E. Diewert. Exact and superlative index numbers. Journal of Econometrics, 4(2):115–145, May 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0304407676900099.

Dubbelman:1978:TCE

[DLA78] C. Dubbelman, A. S. Louter, and A. P. J. Abrahamse. On typical characteristics of economic time series and the relative qualities of five autocorrelation tests. Journal of Econometrics, 8(3): 295–306, December 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407678900490.

Denny:1977:ERV

[DM77] Michael Denny and Doug May. The existence of a real valueadded function in the Canadian manufacturing sector. Journal of Econometrics, 5(1):55–69, January 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900343.

Dent:1978:MCS

[DM78] Warren Dent and An-Sik Min. A Monte Carlo study of autoregressive integrated moving average processes. Journal of Econometrics, 7(1):23–55, February 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407678900040.

Day:1973:CDM

[DN73] Richard H. Day and Jon P. Nelson. A class of dynamic models for describing and projecting industrial development. Journal of Econometrics, 1(2):155–180, June 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440767390016X.

Dreze:1977:BRA

[Drè77] Jacques H. Drèze. Bayesian regression analysis using polyt densities. Journal of Econometrics, 6(3):329–354, November 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407677900045.

Dent:1978:URL

[DS78a] Warren T. Dent and George P. H. Styan. Uncorrelated residuals from linear models. Journal of Econometrics, 7(2):211– 225, June 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900702.

Dwivedi:1978:OLS

[DS78b] T. D. Dwivedi and V. K. Srivastava. Optimality of least squares in the seemingly unrelated regression equation model. Journal of Econometrics, 7(3):391–395, April 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900623.

DeHaan:1978:APC

[DTH78] Laurens De Haan and Elselien Taconis-Haantjes. Asymptotic properties of a correlation coefficient type statistic connected with the general linear model. Journal of Econometrics, 7(1):15–21, February 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900039.

Farebrother:1979:EAD

[Far79] R. W. Farebrother. Estimation with aggregated data. Journal of Econometrics, 10(1):43–55, April 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679900630.

Fuller:1974:ELM

[FB74] Wayne A. Fuller and George E. Battese. Estimation of linear models with crossed-error structure. Journal of Econometrics, 2(1): 67–78, May 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440767490030X.

Fomby:1978:COL

[FG78] Thomas B. Fomby and David K. Guilkey. On choosing the optimal level of significance for the Durbin–Watson test and the Bayesian alternative. Journal of Econometrics, 8(2):203–213, October 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900295.

Ferber:1979:SEE

[FH79] Robert Ferber and Werner Z. Hirsch. Social experiments in economics. Journal of Econometrics, 11(1):77–115, September 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900551.

Farley:1975:SCT

[FHM75] John U. Farley, Melvin Hinich, and Timothy W. McGuire. Some comparisons of tests for a shift in the slopes of a multivariate linear time series model. Journal of Econometrics, 3(3):297– 318, August 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900378.

Fisher:1976:NPE

[Fis76] Walter D. Fisher. Normalization in point estimation. Journal of Econometrics, 4(3):243–252, August 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440767690035X.

Fisher:1978:R

[Fis78] Walter D. Fisher. Reply. Journal of Econometrics, 7(1):127, February 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440767890012X.

Fitts:1973:TAA

[Fit73] John Fitts. Testing for autocorrelation in the autoregressive moving average error model. Journal of Econometrics, 1(4):363–376, December 1973. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407673900225.

Froyen:1974:TEM

[Fro74] Richard T. Froyen. A test of the endogeneity of monetary policy. Journal of Econometrics, 2(2):175–188, July 1974. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407674900384.

Farebrother:1974:GCE

[FS74] R. W. Farebrother and N. E. Savin. The graph of the k-class estimator: An algebraic and statistical interpretation. Journal of Econometrics, 2(4):373–388, December 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407674900219.

Fuss:1977:DEC

[Fus77] Melvyn A. Fuss. The demand for energy in Canadian manufacturing: An example of the estimation of production structures with many inputs. Journal of Econometrics, 5(1):89–116, January 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900367.

Gabrielsen:1978:CI

[Gab78] Arne Gabrielsen. Consistency and identifiability. Journal of Econometrics, 8(2):261–263, October 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900350.

Gallant:1975:SUN

[Gal75] A. Ronald Gallant. Seemingly unrelated nonlinear regressions. Journal of Econometrics, 3(1):35–50, February 1975. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407675900640.

Gallant:1977:TSL

[Gal77] A. Ronald Gallant. Three-stage least-squares estimation for a system of simultaneous, nonlinear, implicit equations. Journal of Econometrics, 5(1):71–88, January 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900355.

Gastwirth:1975:EFM

[Gas75] Joseph L. Gastwirth. The estimation of a family of measures of economic inequality. Journal of Econometrics, 3(1):61–70, February 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900676.

Gaver:1976:BRE

[Gav76] Ken Gaver. Book review: Econometrics and economic theory: Essays in honour of Jan Tinbergen: W. Sellekaerts (International Arts and Sciences Press, White Plains, N.Y., 1974). Journal of Econometrics, 4(4):399, November 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407676900282.

Gill:1978:POA

[GB78] Leonard Gill and Sophocles N. Brissimis. Polynomial operators and the asymptotic distribution of dynamic multipliers. Journal of Econometrics, 7(3):373–384, April 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440767890060X.

Griffiths:1979:BER

[GDP79] William E. Griffiths, Ross G. Drynan, and Surekha Prakash. Bayesian estimation of a random coefficient model. Journal of Econometrics, 10(2):201–220, June 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679900058.

Geraci:1976:ISE

[Ger76] Vincent J. Geraci. Identification of simultaneous equation models with measurement error. Journal of Econometrics, 4(3):263– 283, August 1976. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407676900373.

Granger:1979:RLC

[GERA79] Clive W. J. Granger, Robert Engle, Ramu Ramanathan, and Allan Andersen. Residential load curves and time-of-day pricing: An econometric analysis. Journal of Econometrics, 9(1–2):13– 32, January 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900940.

Geweke:1978:TES

[Gew78] John Geweke. Testing the exogeneity specification in the complete dynamic simultaneous equation model. Journal of Econometrics, 7(2):163–185, June 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407678900672.

Giles:1975:DBA

[Gil75] D. E. A. Giles. Discriminating between autoregressive forms: A Monte Carlo comparison of Bayesian and ad hoc methods. Journal of Econometrics, 3(3):229–248, August 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407675900330.

Gilbert:1977:RUM

[Gil77] Christopher L. Gilbert. Regression using mixed annual and quarterly data. Journal of Econometrics, 5(2):221–239, March 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407677900264.

Gallant:1979:SIS

[GJ79] A. Ronald Gallant and Dale W. Jorgenson. Statistical inference for a system of simultaneous, non-linear, implicit equations in the context of instrumental variable estimation. Journal of Econometrics, 11(2–3):275–302, October/December 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679900411.

Giles:1978:FOA

[GK78] D. E. A. Giles and M. L. King. Fourth-order autocorrelation: Further significance points for the Wallis test. Journal of Econometrics, 8(2):255–259, October 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407678900349.

Gourieroux:1979:CJP

[GM79a] Christian Gourieroux and Alain Monfort. On the characterization of a joint probability distribution by conditional distributions. Journal of Econometrics, 10(1):115–118, April 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679900708.

Gupta:1979:OVV

[GM79b] Yash Pal Gupta and Esfandiar Maasoumi. Omitted variables, variability of estimated parameters and the appearance of autocorrelated disturbances. Journal of Econometrics, 9(3):379–385, February 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900800.

Granger:1974:SRE

[GN74] C. W. J. Granger and P. Newbold. Spurious regressions in econometrics. Journal of Econometrics, 2(2):111–120, July 1974. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407674900347.

Granger:1976:URD

[GN76] C. W. J. Granger and P. Newbold. The use of R^{2} to determine the appropriate transformation of regression variables. Journal of Econometrics, 4(3):205–210, August 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407676900324.

Godfrey:1978:TMH

[God78] Leslie G. Godfrey. Testing for multiplicative heteroskedasticity. Journal of Econometrics, 8(2):227–236, October 1978. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407678900313.

Goldfeld:1973:MMS

[GQ73] Stephen M. Goldfeld and Richard E. Quandt. A Markov model for switching regressions. Journal of Econometrics, 1(1):3–15, March 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440767390002X.

Goldfelfd:1975:EDM

[GQ75] Stephen M. Goldfelfd and Richard E. Quandt. Estimation in a disequilibrium model and the value of information. Journal of Econometrics, 3(4):325–348, November 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407675900524.

Gonedes:1974:BAU

[GR74] Nicholas J. Gonedes and Harry V. Roberts. Bayesian assessment of the unconditional mean square error of repeated predictions from a regression equation. Journal of Econometrics, 2(3):221– 240, September 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900025.

Gonedes:1977:DRW

[GR77] Nicholas J. Gonedes and Harry V. Roberts. Differencing of random walks and near random walks. Journal of Econometrics, 6 (3):289–308, November 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407677900021.

Giles:1979:MSE

[GR79a] D. E. A. Giles and A. C. Rayner. The mean squared errors of the maximum likelihood and natural-conjugate Bayes regression estimators. Journal of Econometrics, 11(2–3):319–334, October/December 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900435.

Gollop:1979:FIO

[GR79b] Frank M. Gollop and Mark J. Roberts. Firm interdependence in oligopolistic markets. Journal of Econometrics, 10(3):313– 331, August 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900873.

Grether:1974:COD

[Gre74] David M. Grether. Correlations with ordinal data. Journal of Econometrics, 2(3):241–246, September 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407674900037.

Grossman:1975:REE

[Gro75] Sanford Grossman. Rational expectations and the econometric modeling of markets subject to uncertainty: A Bayesian approach. Journal of Econometrics, 3(3):255–272, August 1975. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407675900354.

Guilkey:1979:SSS

[GS79] David K. Guilkey and Peter Schmidt. Some small sample properties of estimators and test statistics in the multivariate logit model. Journal of Econometrics, 10(1):33–42, April 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900629.

Guilkey:1974:ATF

[Gui74] David K. Guilkey. Alternative tests for a first-order vector autoregressive error specification. Journal of Econometrics, 2(1): 95–104, May 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900323.

Gunderson:1974:RTS

[Gun74] Morley Gunderson. Retention of trainees: A study with dichotomous dependent variables. Journal of Econometrics, 2(1): 79–93, May 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900311.

Gustafson:1978:TUE

[Gus78] Elizabeth F. Gustafson. Testing unstable econometric models for stability: An empirical study. Journal of Econometrics, 8(2):193– 201, October 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900283.

Guthrie:1976:NBE

[Gut76] Robert S. Guthrie. A note on the Bayesian estimation of Solow’s distributed lag model. Journal of Econometrics, 4(3):295–300, August 1976. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407676900397.

Goldberg:1977:SAP

[GV77] Michael A. Goldberg and Ashok Vora. Spectral analysis of public utility returns. Journal of Econometrics, 6(1):79–101, July 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407677900562.

Haddad:1976:TIS

[Had76] C. L. Haddad. Testing income series: An application of principal components. Journal of Econometrics, 4(1):27–40, February 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407676900154.

Hasenkamp:1976:SMO

[Has76] Georg Hasenkamp. A study of multiple-output production functions: Klein’s railroad study revisited. Journal of Econometrics, 4(3):253–262, August 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407676900361.

Hatanaka:1974:ETS

[Hat74] Michio Hatanaka. An efficient two-step estimator for the dynamic adjustment model with autoregressive errors. Journal of Econometrics, 2(3):199–220, September 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407674900013.

Hatanaka:1976:SET

[Hat76] Michio Hatanaka. Several efficient two-step estimators for the dynamic simultaneous equations model with autoregressive disturbances. Journal of Econometrics, 4(2):189–204, May 1976. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0304407676900129.

Hatanaka:1978:EEM

[Hat78] Michio Hatanaka. On the efficient estimation methods for the macro-economic models nonlinear in variables. Journal of Econometrics, 8(3):323–356, December 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900519.

Hausman:1977:EVS

[Hau77] Jerry A. Hausman. Errors in variables in simultaneous equation models. Journal of Econometrics, 5(3):389–401, May 1977. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407677900471.

Harvey:1977:TFM

[HC77] Andrew C. Harvey and Patrick Collier. Testing for functional misspecification in regression analysis. Journal of Econometrics, 6(1): 103–119, July 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900574.

Hendry:1976:SSE

[Hen76] David F. Hendry. The structure of simultaneous equations estimators. Journal of Econometrics, 4(1):51–88, February 1976. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407676900178.

Hendry:1979:BII

[Hen79] David F. Hendry. The behaviour of inconsistent instrumental variables estimators in dynamic systems with autocorrelated errors. Journal of Econometrics, 9(3):295–314, February 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900769.

Higgins:1973:EMA

[HF73] C. I. Higgins and V. W. Fitzgerald. An econometric model of the Australian economy. Journal of Econometrics, 1(3):229– 265, October 1973. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407673900092.

Hendry:1974:MCM

[HH74] David F. Hendry and Robin W. Harrison. Monte Carlo methodology and the small sample behaviour of ordinary and twostage least squares. Journal of Econometrics, 2(2):151–174, July 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407674900372.

Hamlen:1978:HAA

[HH78] Susan S. Hamlen and William A. Hamlen. Harmonic alternatives to the Almon polynomial technique. Journal of Econometrics, 7 (1):57–66, February 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407678900052.

Hausmann:1979:TLE

[HKM79] J. A. Hausmann, M. Kinnucan, and D. McFaddden. A twolevel electricity demand model: Evaluation of the Connecticut time-of-day pricing test. Journal of Econometrics, 10(3):263– 289, August 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440767990085X.

Hendricks:1977:CPE

[HKP77] Wallace Hendricks, Roger Koenker, and Robert Podlasek. Consumption patterns for electricity. Journal of Econometrics, 5(2): 135–153, March 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900203.

Hendricks:1979:RDE

[HKP79] Wallace Hendricks, Roger Koenker, and Dale J. Poirier. Residential demand for electricity: An econometric approach. Journal of Econometrics, 9(1–2):33–57, January 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679900952.

Ho:1979:RID

[Ho79] Dit Sang Ho. The reconstruction of index data in aggregative econometric models. Journal of Econometrics, 10(3):333– 359, August 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900885.

Holland:1977:MSE

[Hol77] Burt S. Holland. Mean square error tests for restrictions in singular linear models. Journal of Econometrics, 6(3):355–363, November 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900057.

Harvey:1974:CPS

[HP74] A. C. Harvey and G. D. A. Phillips. A comparison of the power of some tests for heteroskedasticity in the general linear model. Journal of Econometrics, 2(4):307–316, December 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407674900165.

Hartley:1974:EPL

[HR74] Michael J. Hartley and Nagesh S. Revankar. On the estimation of the Pareto law from under-reported data. Journal of Econometrics, 2(4):327–341, December 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407674900189.

Hinkley:1977:EPL

[HR77] David V. Hinkley and Nagesh S. Revankar. Estimation of the Pareto law from underreported data: A further analysis. Journal of Econometrics, 5(1):1–11, January 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900318.

Hymans:1976:AHI

[HS76] Saul H. Hymans and Harold T. Shapiro. The allocation of household income to food consumption. Journal of Econometrics, 4(2): 167–188, May 1976. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0304407676900117.

Hsiao:1979:LRU

[Hsi79] Cheng Hsiao. Linear regression using both temporally aggregated and temporally disaggregated data. Journal of Econometrics, 10(2): 243–252, June 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900083.

Huntzinger:1979:MAR

[Hun79] R. La Var Huntzinger. Market analysis with rational expectations: Theory and estimation. Journal of Econometrics, 10(2): 127–145, June 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900010.

Hurd:1973:BRI

[Hur73] Michael D. Hurd. Book review: An introduction to applied macroeconomics: Edwin Kuh and Richard Schmalensee, (North-Holland Publ. Co., Amsterdam, 1972) 240 pp. (Dfl. 40.00). Journal of Econometrics, 1(4):403–406, December 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407673900304.

Hurd:1979:ETS

[Hur79] Michael Hurd. Estimation in truncated samples when there is heteroscedasticity. Journal of Econometrics, 11(2–3):247–258, October/December 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900393.

Harkema:1977:BNB

[HV77] Rins Harkema and Sybrand Schim Van Der Loeff. On Bayesian and non-Bayesian estimation of a two-level CES production function for the Dutch manufacturing sector. Journal of Econometrics, 5(2):155–165, March 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407677900215.

Hylleberg:1977:CSF

[Hyl77] Svend Hylleberg. A comparative study of finite sample properties of band spectrum regression estimators. Journal of Econometrics, 5(2):167–182, March 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407677900227.

Just:1978:SSP

[JP78] Richard E. Just and Rulon D. Pope. Stochastic specification of production functions and economic implications. Journal of Econometrics, 7(1):67–86, February 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407678900064.

Judge:1975:BRA

[Jud75] George Judge. Book review: Applied multivariate analysis: S. James Press, (Holt, Rinehart and Winston, New York, 1972) xix + 521 pp. Journal of Econometrics, 3(3):320, August 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407675900391.

Judge:1973:PEA

[JYB73] G. G. Judge, T. A. Yancey, and M. E. Bock. Properties of estimators after preliminary tests of significance when stochastic restrictions are used in regression. Journal of Econometrics, 1(1): 29–47, March 1973. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407673900043.

Kadane:1976:BRS

[Kad76] Joseph B. Kadane. Book review: Structural equation model in the social sciences: A. S. Goldberger and O. D. Duncan, eds. (Seminar Press, New York, 1973). Journal of Econometrics, 4(1): 97, February 1976. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407676900191.

Kadane:1978:CNP

[Kad78a] Joseph B. Kadane. A comment on “normalization in point estimation”. Journal of Econometrics, 7(1):123–125, February 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407678900118.

Kadane:1978:R

[Kad78b] Joseph B. Kadane. Rejoinder. Journal of Econometrics, 7(1): 129, February 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900131.

Kakwani:1977:EEE

[Kak77] Nanak Kakwani. On the estimation of Engel elasticities from grouped observations with application to Indonesian data. Journal of Econometrics, 6(1):1–19, July 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900513.

Kakwani:1978:NME

[Kak78] Nanak Kakwani. A new method of estimating Engel elasticities. Journal of Econometrics, 8(1):103–110, August 1978. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407678900921.

Kassouf:1976:LSO

[Kas76] S. T. Kassouf. The lag structure of option price. Journal of Econometrics, 4(4):303–310, November 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440767690021X.

Keller:1975:NCL

[Kel75] Wouter J. Keller. A new class of limited-information estimators for simultaneous equations systems. Journal of Econometrics, 3 (1):71–92, February 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407675900688.

Kenward:1975:ADM

[Ken75] Lloyd R. Kenward. Autocorrelation and dynamic methodology with an application to wage determination models. Journal of Econometrics, 3(2):179–187, May 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407675900469.

Kiefer:1978:FSO

[Kie78] Nicholas M. Kiefer. Federally subsidized occupational training and the employment and earnings of male trainees. Journal of Econometrics, 8(1):111–125, August 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900933.

Kraft:1974:EEV

[KK74] John Kraft and Arthur Kraft. Empirical estimation of the value of travel time using multi mode choice models. Journal of Econometrics, 2(4):317–326, December 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407674900177.

Klein:1979:OIW

[Kle79a] Roger W. Klein. Optimal instruments when the disturbances are small. Journal of Econometrics, 9(3):368–377, February 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900794.

Klevmarken:1979:CSC

[Kle79b] N. Anders Klevmarken. A comparative study of complete systems of demand functions. Journal of Econometrics, 10(2):165– 191, June 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900034.

Knight:1977:EMP

[Kni77] John L. Knight. On the existence of moments of the partially restricted reduced-form estimators from a simultaneousequation model. Journal of Econometrics, 5(3):315–321, May 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407677900422.

Koenker:1979:OPL

[Koe79] Roger Koenker. Optimal peak load pricing with time-additive consumer preferences. Journal of Econometrics, 9(1–2):175–192, January 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679901027.

Kohn:1978:LGI

[Koh78] R. Kohn. Local and global identification and strong consistency in time series models. Journal of Econometrics, 8(3):269–293, December 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900489.

Kloek:1978:EEI

[KvD78] Teun Kloek and Herman K. van Dijk. Efficient estimation of income distribution parameters. Journal of Econometrics, 8(1):61– 74, August 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900908.

Laffargue:1977:NMA

[Laf77] Jean-Pierre Laffargue. Nonlinear models of analysis of variance. Journal of Econometrics, 5(3):347–363, May 1977. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407677900446.

Lancaster:1979:PBC

[Lan79] Tony Lancaster. Prediction from binary choice models: A note. Journal of Econometrics, 9(3):387–389, February 1979. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407679900812.

Lapp:1976:BRB

[Lap76] John S. Lapp. Book review: Bank management and portfolio behavior: Donald D. Hester and James L. Pierce (Yale University Press, New Haven, 1975). Journal of Econometrics, 4(1):97– 99, February 1976. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407676900208.

Lawrence:1979:RDE

[LB79] Anthony Lawrence and Steven Braithwait. The residential demand for electricity with time-of-day pricing. Journal of Econometrics, 9(1–2):59–77, January 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407679900964.

Leamer:1975:RSR

[Lea75] Edward E. Leamer. A result on the sign of restricted leastsquares estimates. Journal of Econometrics, 3(4):387–390, November 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440767590055X.

Lee:1973:BRN

[Lee73] T. C. Lee. Book review: Nonlinear methods in econometrics: S. M. Goldfeld and R. E. Quandt, (North-Holland Publ. Co., Amsterdam and London, 1972) xii + 280 pp. ($18.75). Journal of Econometrics, 1(4):399–401, December 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407673900274.

Levi:1977:BRE

[Lev77a] Maurice D. Levi. Book review: Econometrics of investment: J. C. R. Rowley and P. K. Trivedi, (Wiley, New York, 1975). Journal of Econometrics, 6(1):142, July 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900616.

Levi:1977:MEB

[Lev77b] Maurice D. Levi. Measurement errors and bounded OLS estimates. Journal of Econometrics, 6(2):165–171, September 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407677900136.

Laffont:1979:DED

[LM79] Jean-Jacques Laffont and Alain Monfort. Disequilibrium econometrics in dynamic models. Journal of Econometrics, 11(2–3):353– 361, October/December 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407679900459.

Lovell:1973:NAB

[Lov73] C. A. Knox Lovell. A note on aggregation bias and loss. Journal of Econometrics, 1(3):301–311, October 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407673900110.

LEsperance:1975:PFT

[LT75] Wilford L. L’Esperance and Daniel Taylor. The power of four tests of autocorrelation in the linear regression model. Journal of Econometrics, 3(1):1–21, February 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407675900627.

Lee:1978:ESL

[LT78a] Lung-Fei Lee and Robert P. Trost. Estimation of some limited dependent variable models with application to housing demand. Journal of Econometrics, 8(3):357–382, December 1978. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407678900520.

Lee:1978:SFP

[LT78b] Lung-Fei Lee and William G. Tyler. The stochastic frontier production function and average efficiency: An empirical analysis. Journal of Econometrics, 7(3):385–389, April 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900611.

Maddala:1976:BRS

[Mad76] G. S. Maddala. Book review: Studies in Bayesian Econometrics and Statistics, In honor of Leonard J. Savage: S. E. Fienberg and A. Zellner (North-Holland Publishing Co., Amsterdam, 1975). Journal of Econometrics, 4(4):399–400, November 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407676900294.

Magnus:1978:MLE

[Mag78] Jan R. Magnus. Maximum likelihood estimation of the GLS model with unknown parameters in the disturbance covariance matrix. Journal of Econometrics, 7(3):281–312, April 1978. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407678900568.

Maital:1978:MSS

[Mai78] Shlomo Maital. Multidimensional scaling: Some econometric applications. Journal of Econometrics, 8(1):33–46, August 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440767890088X.

Manski:1975:MSE

[Man75] Charles F. Manski. Maximum score estimation of the stochastic utility model of choice. Journal of Econometrics, 3(3):205– 228, August 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900329.

Mariano:1975:SLC

[Mar75] Roberto S. Mariano. Some large-concentration-parameter asymptotics for the k-class estimators. Journal of Econometrics, 3(2): 171–177, May 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900457.

Maravall:1976:NTS

[Mar76] Agustin Maravall. A note on three-stage least squares estimation. Journal of Econometrics, 4(4):325–330, November 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407676900233.

Mazodier:1975:BRE

[Maz75] Pascal Mazodier. Book review: Efficient estimation with a priori information, Cowless Foundation Monograph no. 23: Thomas J. Rothenberg, (Yale university press, New York, 1973) viii + 180 pp. $10.00. Journal of Econometrics, 3(1):99–102, February 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407675900755.

McCallum:1974:RIM

[McC74] B. T. McCallum. The relative impact of monetary and fiscal policy instruments: Some structure-based estimates. Journal of Econometrics, 2(3):283–299, September 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407674900062.

McElroy:1977:GFS

[McE77a] Marjorie B. McElroy. Goodness of fit for seemingly unrelated regressions: Glahn’s R_{y.x}^{2} and Hooper’s r̄^{2}. Journal of Econometrics, 6(3):381–387, November 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407677900082.

McElroy:1977:WMC

[McE77b] Marjorie B. McElroy. Weaker MSE criteria and tests for linear restrictions in regression models with non-spherical disturbances. Journal of Econometrics, 6(3):389–394, November 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407677900094.

Mentz:1977:EFO

[Men77] Raúl Pedro Mentz. Estimation in the first-order moving average model through the finite autoregressive approximation: Some asymptotic results. Journal of Econometrics, 6(2):225–236, September 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407677900161.

Mitchell:1977:BRE

[Mit77] Bridger M. Mitchell. Book review: Econometric studies of U.S. energy policy: D. W. Jorgenson, ed., (North-Holland, Amsterdam, 1976) pp. 243. Journal of Econometrics, 6(1):141– 142, July 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900604.

Manning:1979:DAP

[MMA79] Williard G. Manning, Bridger M. Mitchell, and Jan Paul Acton. Design of the Los angeles peak-load pricing experiment for electricity. Journal of Econometrics, 11(1):131–194, September 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900575.

Mehta:1978:EDD

[MNS78] Jatinder S. Mehta, Gorti V. L. Narasimham, and Paravastu A. V. B. Swamy. Estimation of a dynamic demand function for gasoline with different schemes of parameter variation. Journal of Econometrics, 7(3):263–279, April 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407678900556.

Monfort:1978:FOI

[Mon78] Alain Monfort. First-order identification in linear models. Journal of Econometrics, 7(3):333–350, April 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900581.

Morris:1979:FSM

[Mor79] Carl Morris. A finite selection model for experimental design of the health insurance study. Journal of Econometrics, 11(1):43– 61, September 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900538.

McMenamin:1978:SED

[MP78] J. Stuart McMenamin and Jean-Paul Pinard. Specification and estimation of dynamic demand systems incorporating polynomial price response functions: An application to U.S. clothing imports. Journal of Econometrics, 7(2):147–162, June 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900660.

Mehta:1978:EMS

[MS78] Jatinder S. Mehta and Paravastu A. V. B. Swamy. The existence of moments of some simple Bayes estimators of coefficients in a simultaneous equation model. Journal of Econometrics, 7(1):1– 13, February 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900027.

Chaudhuri:1977:ADL

[(Mu77] Maitreyi Chaudhuri (Mukherjee). Autocorrelated disturbances in the light of specification analysis. Journal of Econometrics, 5(3): 301–313, May 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900410.

Morgan:1974:THS

[MV74] Alison Morgan and Waiter Vandaele. On testing hypothesis in simultaneous equation models. Journal of Econometrics, 2(1): 55–65, May 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900293.

Neftci:1979:MPS

[Nef79] Salih N. Neftçi. Modeling the price side of econometric models: An analysis of the underlying hypotheses. Journal of Econometrics, 10(1):71–84, April 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407679900654.

Nelson:1974:FOM

[Nel74] Charles R. Nelson. The first-order moving average process: Identification, estimation and prediction. Journal of Econometrics, 2(2): 121–141, July 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900359.

Nelson:1976:GEJ

[Nel76] Charles R. Nelson. Gains in efficiency from joint estimation of systems of autoregressive-moving average processes. Journal of Econometrics, 4(4):331–348, November 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407676900245.

Nelson:1977:CRM

[Nel77] Forrest D. Nelson. Censored regression models with unobserved, stochastic censoring thresholds. Journal of Econometrics, 6(3): 309–327, November 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407677900033.

Neudecker:1977:AKN

[Neu77] Heinz Neudecker. Abrahamse and Koerts’ ‘new estimator’ of disturbances in regression analysis. Journal of Econometrics, 5(1):129– 133, January 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900380.

Nelson:1979:EUB

[NG79] Harold L. Nelson and C. W. J. Granger. Experience with using the Box–Cox transformation when forecasting economic time series. Journal of Econometrics, 10(1):57–69, April 1979. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407679900642.

Newhouse:1979:MIS

[NMM^{+}79] Joseph P. Newhouse, Kent H. Marquis, Carl N. Morris, Charles E. Phelps, and William H. Rogers. Measurement issues in the second generation of social experiments: The health insurance study. Journal of Econometrics, 11(1):117–129, September 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900563.

Nakamura:1978:ITS

[NN78] Alice Nakamura and Masao Nakamura. On the impact of the tests for serial correlation upon the test of significance for the regression coefficient. Journal of Econometrics, 7(2):199– 210, June 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900696.

Nagar:1978:BMS

[NS78] Anirudh L. Nagar and Surottam N. Sahay. The bias and mean squared error of forecasts from partially restricted reduced form. Journal of Econometrics, 7(2):227–243, June 1978. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407678900714.

Nelson:1979:HTB

[NS79] Charles R. Nelson and Gary S. Shea. Hypothesis testing based on goodness-of-fit in the moving average time series model. Journal of Econometrics, 10(2):221–226, June 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440767990006X.

OBrien:1975:BRQ

[O’B75] R. J. O’Brien. Book review: Quantitative methods in Economics: J. D. A. Cuddy, (Rotterdam University Press, 1974) pp. viii + 180, Dfl. 37.50 (U.S. $13.20). Journal of Econometrics, 3(2): 200–201, May 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900494.

OHara:1977:BRT

[O’H77] Donald O’Hara. Book review: The theory of quantitative economic policy: K. A. Fox, J. K. Sengupta and E. Thorbecke, 2nd rev. ed. (North-Holland, Amsterdam, 1973). Journal of Econometrics, 5(2):259, March 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407677900288.

Pagan:1973:BRE

[Pag73a] Adrian Pagan. Book review: Econometric studies of macro and monetary relations: A. A. Powell and R. A. Williams (eds.), (North-Holland Publ. Co., Amsterdam, 1973) viii + 358 pp. ($18.75). Journal of Econometrics, 1(4):402–403, December 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407673900298.

Pagan:1973:EEM

[Pag73b] Adrian Pagan. Efficient estimation of models with composite disturbance terms. Journal of Econometrics, 1(4):329–340, December 1973. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407673900195.

Pagan:1978:RPL

[Pag78] Adrian Pagan. Rational and polynomial lags: The finite connection. Journal of Econometrics, 8(2):247–254, October 1978. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407678900337.

Palm:1977:UTS

[Pal77] Franz Palm. On univariate time series methods and simultaneous equation econometric models. Journal of Econometrics, 5(3): 379–388, May 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440767790046X.

Panton:1976:CBC

[Pan76] D. Panton. Chicago board call options as predictors of common stock price changes. Journal of Econometrics, 4(2):101–113, May 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0304407676900087.

Papakyriazis:1978:OED

[Pap78] Panagiotis A. Papakyriazis. Optimal experimental design in econometrics: The time series problem. Journal of Econometrics, 7(3): 351–372, April 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900593.

Pierce:1977:CTS

[PH77] David A. Pierce and Larry D. Haugh. Causality in temporal systems: Characterization and a survey. Journal of Econometrics, 5(3):265–293, May 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407677900392.

Pierce:1979:CIC

[PH79] David A. Pierce and Larry D. Haugh. The characterization of instantaneous causality: A comment. Journal of Econometrics, 10(2):257–259, June 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407679900101.

Phillips:1973:PIF

[Phi73] P. C. B. Phillips. The problem of identification in finite parameter continuous time models. Journal of Econometrics, 1(4):351– 362, December 1973. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407673900213.

Phillips:1977:RTS

[Phi77a] Garry D. A. Phillips. Recursions for the two-stage leastsquares estimators. Journal of Econometrics, 6(1):65–77, July 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407677900550.

Phillips:1977:AFS

[Phi77b] Peter C. B. Phillips. An approximation to the finite sample distribution of Zellner’s seemingly unrelated regression estimator. Journal of Econometrics, 6(2):147–164, September 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900124.

Phillips:1979:CER

[Phi79a] P. C. B. Phillips. The concentration ellipsoid of a random vector. Journal of Econometrics, 11(2–3):363–365, October/ December 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900460.

Phillips:1979:SDF

[Phi79b] Peter C. B. Phillips. The sampling distribution of forecasts from a first-order autoregression. Journal of Econometrics, 9(3):241– 261, February 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900733.

Pierce:1975:FDM

[Pie75] David A. Pierce. Forecasting in dynamic models with stochastic regressors. Journal of Econometrics, 3(4):349–374, November 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407675900536.

Plosser:1979:ASE

[Plo79] Charles I. Plosser. The analysis of seasonal economic models. Journal of Econometrics, 10(2):147–163, June 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679900022.

Poirier:1975:BRR

[Poi75a] Dale J. Poirier. Book review: Regression estimation from grouped observations, Griffin’s statistical monographs and courses no. 33: Y. Haitovosky, (Griffin and co., London) x + 94 pp., £2.30. Journal of Econometrics, 3(1):93, February 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440767590069X.

Poirier:1975:UBS

[Poi75b] Dale J. Poirier. On the use of bilinear splines in economics. Journal of Econometrics, 3(1):23–34, February 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407675900639.

Parzen:1979:AMS

[PP79] Emanuel Parzen and Marcello Pagano. An approach to modeling seasonally stationary time series. Journal of Econometrics, 9(1– 2):137–153, January 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407679901003.

Price:1979:CIC

[Pri79] J. Michael Price. The characterization of instantaneous causality: A correction. Journal of Econometrics, 10(2):253–256, June 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900095.

Plosser:1977:ENI

[PS77] Charles I. Plosser and G. William Schwert. Estimation of a non-invertible moving average process: The case of overdifferencing. Journal of Econometrics, 6(2):199–224, September 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440767790015X.

Panzar:1979:TDI

[PW79] John C. Panzar and Robert D. Willig. Theoretical determinants of the industrial demand for electricity by time of day. Journal of Econometrics, 9(1–2):193–207, January 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679901039.

Press:1978:PDM

[PZ78] S. James Press and Arnold Zellner. Posterior distribution for the multiple correlation coefficient with fixed regressors. Journal of Econometrics, 8(3):307–321, December 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900507.

Raduchel:1978:NNL

[Rad78] William J. Raduchel. A note on non-linear limited-information maximum-likelihood. Journal of Econometrics, 7(1):119–122, February 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900106.

Reinsel:1979:FED

[Rei79] Greg Reinsel. FIML estimation of the dynamic simultaneous equations model with ARMA disturbances. Journal of Econometrics, 9 (3):263–281, February 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407679900745.

Richard:1975:NIM

[Ric75a] Jean-François Richard. A note on the information matrix of the multivariate normal distribution. Journal of Econometrics, 3(1): 57–60, February 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900664.

Richardson:1975:BRI

[Ric75b] J. David Richardson. Book review: The international linkage of national economic models: R. J. Ball, ed., (North-Holland, Amsterdam, 1973) xii + pp., $31.50. Journal of Econometrics, 3 (1):96–97, February 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407675900731.

Robinson:1977:CEC

[Rob77] Peter M. Robinson. The construction and estimation of continuous time models and discrete approximations in econometrics. Journal of Econometrics, 6(2):173–197, September 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900148.

Rose:1977:FAI

[Ros77] David E. Rose. Forecasting aggregates of independent Arima processes. Journal of Econometrics, 5(3):323–345, May 1977. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407677900434.

Rice:1977:EMP

[RS77] Patricia Rice and V. Kerry Smith. An econometric model of the petroleum industry. Journal of Econometrics, 6(3):263–287, November 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440767790001X.

Salkever:1976:UDV

[Sal76] David S. Salkever. The use of dummy variables to compute predictions, prediction errors, and confidence intervals. Journal of Econometrics, 4(4):393–397, November 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407676900270.

Sasaki:1978:EAL

[Sas78] Komei Sasaki. An empirical analysis of linear aggregation problems: The case of investment behavior in Japanese firms. Journal of Econometrics, 7(3):313–331, April 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440767890057X.

Sawa:1973:MSE

[Saw73] Takamitsu Sawa. The mean square error of a combined estimator and numerical comparison with the TSLS estimator. Journal of Econometrics, 1(2):115–132, June 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407673900146.

Sawa:1978:EML

[Saw78] Takamitsu Sawa. The exact moments of the least squares estimator for the autoregressive model. Journal of Econometrics, 8 (2):159–172, October 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407678900258.

Spann:1979:EEP

[SB79] Robert M. Spann and Edward C. Beauvais. Econometric estimation of peak electricity demands. Journal of Econometrics, 9(1– 2):119–136, January 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/030440767990099X.

Schonfeld:1973:NMP

[Sch73] Peter Schönfeld. A note on the measurability of the pseudoinverse. Journal of Econometrics, 1(3):313–314, October 1973. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407673900122.

Schonfeld:1975:BRM

[Sch75a] P. Schönfeld. Book review: Multicollinearity in linear economic models, Tilburg studies in economics no. 7: D. Neelman, (Tilburg university press, The Netherlands, 1973) viii + 103 pp. Journal of Econometrics, 3(1):94–95, February 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407675900718.

Schonfeld:1975:NLS

[Sch75b] Peter Schönfeld. A note on least squares estimation and the blue in a generalized linear regression model. Journal of Econometrics, 3(2):189–197, May 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407675900470.

Scharf:1976:MCE

[Sch76] Werner Scharf. K-matrix-class estimators and the full-information maximum-likelihood estimator as a special case. Journal of Econometrics, 4(1):41–50, February 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407676900166.

Schmidt:1977:ESU

[Sch77] Peter Schmidt. Estimation of seemingly unrelated regressions with unequal numbers of observations. Journal of Econometrics, 5(3): 365–377, May 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900458.

Schmidt:1978:NES

[Sch78] Peter Schmidt. A note on the estimation of seemingly unrelated regression systems. Journal of Econometrics, 7(2):259– 261, June 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900738.

Srivastava:1979:ESU

[SD79] V. K. Srivastava and T. D. Dwivedi. Estimation of seemingly unrelated regression equations: A brief survey. Journal of Econometrics, 10(1):15–32, April 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407679900617.

Schmidt:1976:EVT

[SG76] Peter Schmidt and David K. Guilkey. The effects of various treatments of truncation remainders on tests of hypotheses in distributed lag models. Journal of Econometrics, 4(3):211– 230, August 1976. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407676900336.

Scobie:1975:EES

[SJ75] Grant M. Scobie and Paul R. Johnson. Estimation of the elasticity of substitution in the presence of errors of measurement. Journal of Econometrics, 3(1):51–56, February 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407675900652.

Schmidt:1979:ETA

[SL79] Peter Schmidt and C. A. Knox Lovell. Estimating technical and allocative inefficiency relative to stochastic production and cost frontiers. Journal of Econometrics, 9(3):343–366, February 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900782.

Swamy:1975:BES

[SM75] P. A. V. B. Swamy and J. S. Mehta. On Bayesian estimation of seemingly unrelated regressions when some observations are missing. Journal of Econometrics, 3(2):157–169, May 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407675900445.

Swamy:1979:ECC

[SM79] P. A. V. B. Swamy and J. S. Mehta. Estimation of common coefficients in two regression equations. Journal of Econometrics, 10 (1):1–14, April 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900605.

Sowey:1973:CBM

[Sow73] Eric R. Sowey. A classified bibliography of Monte Carlo studies in econometrics. Journal of Econometrics, 1(4):377–395, December 1973. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407673900237.

Spencer:1975:SSB

[Spe75] Byron G. Spencer. The small sample bias of Durbin’s tests for serial correlation: When one of the regressors is the lagged dependent variable and the null hypothesis is true. Journal of Econometrics, 3(3):249–254, August 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407675900342.

Spencer:1979:EDS

[Spe79] David E. Spencer. Estimation of a dynamic system of seemingly unrelated regressions with autoregressive disturbances. Journal of Econometrics, 10(2):227–241, June 1979. CODEN JECMB6.

ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407679900071.

Spitzer:1977:SES

[Spi77] John J. Spitzer. A simultaneous equations system of money demand and supply using generalized functional forms. Journal of Econometrics, 5(1):117–128, January 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407677900379.

Swamy:1975:RES

[SR75] P. A. V. B. Swamy and Paul N. Rappoport. Relative efficiencies of some simple Bayes estimators of coefficients in dynamic models — I. Journal of Econometrics, 3(3):273–296, August 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407675900366.

Swamy:1978:RES

[SR78] Paravastu A. V. B. Swamy and Paul N. Rappoport. Relative efficiencies of some simple Bayes estimators of coefficients in a dynamic equation with serially correlated errors — II. Journal of Econometrics, 7(2):245–258, June 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407678900726.

Srivastava:1973:EIM

[Sri73] V. K. Srivastava. The efficiency of an improved method of estimating seemingly unrelated regression equations. Journal of Econometrics, 1(4):341–350, December 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407673900201.

Stafford:1974:BRP

[Sta74] Frank P. Stafford. Book review: Permanent income, wealth, and consumption: A critique of the permanent income theory, the life cycle hypothesis, and related theories: T. Mayer (University of California Press, Berkeley, 1973). Journal of Econometrics, 2(2): 195–196, July 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900402.

Stigum:1976:LSS

[Sti76] Bernt P. Stigum. Least squares and stochastic difference equations. Journal of Econometrics, 4(4):349–370, November 1976. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407676900257.

Savin:1978:ETF

[SW78] N. E. Savin and Kenneth J. White. Estimation and testing for functional form and autocorrelation: A simultaneous approach. Journal of Econometrics, 8(1):1–12, August 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900854.

Szroeter:1978:GVR

[Szr78] Jerzy Szroeter. Generalized variance-ratio tests for serial correlation in multivariate regression models. Journal of Econometrics, 8(1):47–59, August 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407678900891.

Taga:1975:BRS

[Tag75] Yasushi Taga. Book review: Statistical theory of sample survey design and analysis: H. S. Konijn, (North-Holland, Amsterdam, 1974) pp.xv + 429, $32.50. Journal of Econometrics, 3(2): 201–202, May 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900500.

Takayama:1973:BRS

[Tak73] T. Takayama. Book review: Stochastic programming — Methods and applications: Jati K. Sengupta, (North-Holland Publ. Co., Amsterdam, 1973) xi + 313 pp. ($10.00). Journal of Econometrics, 1 (4):398–399, December 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407673900262.

Taub:1979:PCV

[Tau79] Allan J. Taub. Prediction in the context of the variancecomponents model. Journal of Econometrics, 10(1):103–107, April 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/030440767990068X.

Taylor:1979:MRD

[Tay79] Lester D. Taylor. On modelling the residential demand for electricity by time-of-day. Journal of Econometrics, 9(1–2):97–115, January 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679900988.

Trivedi:1973:RII

[Tri73] P. K. Trivedi. Retail inventory investment behaviour. Journal of Econometrics, 1(1):61–80, March 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407673900067.

Trivedi:1975:BRM

[Tri75] P. K. Trivedi. Book review: Macroeconomic regulation: K. P. Vishwakarma, (Rotterdam University Press, 1974) pp. xi + 91. Journal of Econometrics, 3(2):199–200, May 1975. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407675900482.

Tsurumi:1976:BTP

[Tsu76] Hiroki Tsurumi. A Bayesian test of the product cycle hypothesis applied to Japanese crude steel production. Journal of Econometrics, 4(4):371–392, November 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407676900269.

Tsurumi:1977:BTP

[Tsu77] Hiroki Tsurumi. A Bayesian test of a parameter shift and an application. Journal of Econometrics, 6(3):371–380, November 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407677900070.

Tsurumi:1976:BEM

[TT76] Hiroki Tsurumi and Yoshi Tsurumi. A Bayesian estimation of macro and micro CES production functions. Journal of Econometrics, 4(1):1–25, February 1976. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407676900142.

Tiao:1977:SAB

[TTC77] George C. Tiao, Wei-Yuan Tan, and Yu-Chi Chang. Some aspects of bivariate regression subject to linear constraints. Journal of Econometrics, 5(1):13–35, January 1977. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440767790032X.

Toyoda:1975:EVA

[TW75] T. Toyoda and T. D. Wallace. Estimation of variance after a preliminary test of homogeneity and optimal levels of significance for the pre-test. Journal of Econometrics, 3(4):395–404, November 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407675900573.

Toyoda:1979:PTP

[TW79] T. Toyoda and T. Dudley Wallace. Pre-testing on part of the data. Journal of Econometrics, 10(1):119–123, April 1979. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030440767990071X.

Tishler:1979:SRM

[TZ79] Asher Tishler and Israel Zang. A switching regression method using inequality conditions. Journal of Econometrics, 11(2–3):259– 274, October/December 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/030440767990040X.

Ullah:1974:SDI

[Ull74] Aman Ullah. On the sampling distribution of improved estimators for coefficients in linear regression. Journal of Econometrics, 2(2):143–150, July 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407674900360.

Uri:1978:DCM

[UM78] Noel D. Uri and J. Wilson Mixon. The distribution of changes in manufacturing employment and the impact of the minimum wage. Journal of Econometrics, 7(1):103–114, February 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407678900088.

Uri:1979:MTS

[Uri79] Noel D. Uri. A mixed time-series/econometric approach to forecasting peak system load. Journal of Econometrics, 9(1–2):155– 171, January 1979. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407679901015.

Vinod:1976:CRE

[Vin76] H. D. Vinod. Canonical ridge and econometrics of joint production. Journal of Econometrics, 4(2):147–166, May 1976. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0304407676900105.

Visco:1978:ORS

[Vis78] Ignazio Visco. On obtaining the right sign of a coefficient estimate by omitting a variable from the regression. Journal of Econometrics, 7(1):115–117, February 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/030440767890009X.

Wales:1977:FFF

[Wal77] Terence J. Wales. On the flexibility of flexible functional forms: An empirical approach *. Journal of Econometrics, 5(2):183– 193, March 1977. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407677900239.

Wales:1978:LSC

[Wal78] Terence J. Wales. Labour supply and commuting time: An empirical study. Journal of Econometrics, 8(2):215–226, October 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407678900301.

Weiss:1975:BRS

[Wei75] Leonard W. Weiss. Book review: Supply relationships in the Canadian economy, international business and economics studies: Lawrence H. Officer, (Michigan state university, East Lansing, 1972) x + 173 pp. Journal of Econometrics, 3(1):95–96, February 1975. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440767590072X.

Wei:1978:ETA

[Wei78] William W. S. Wei. The effect of temporal aggregation on parameter estimation in distributed lag model. Journal of Econometrics, 8 (2):237–246, October 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0304407678900325.

Westin:1974:PBC

[Wes74] Richard B. Westin. Predictions from binary choice models. Journal of Econometrics, 2(1):1–16, May 1974. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440767490027X.

Westin:1978:PLT

[WG78] Richard B. Westin and David W. Gillen. Parking location and transit demand: A case study of endogenous attributes in disaggregate mode choice models. Journal of Econometrics, 8(1):75– 101, August 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/030440767890091X.

Williamson:1974:BRA

[Wil74] Jeffrey G. Williamson. Book review: Analysis of development problems: Studies of the Chilean economy: R. S. Eckhaus and P. N. Rosentein-Rodan, eds., (North-Holland, Amsterdam, 1973) xii + 430 pp., $30.00. Journal of Econometrics, 2(4):393–394, December 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900232.

Wills:1979:TCU

[Wil79] John Wills. Technical change in the U.S. primary metals industry. Journal of Econometrics, 10(1):85–98, April 1979. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407679900666.

Woodland:1978:TWS

[Woo78] Alan D. Woodland. On testing weak separability. Journal of Econometrics, 8(3):383–398, December 1978. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0304407678900532.

Woodland:1979:SSE

[Woo79] A. D. Woodland. Stochastic specification and the estimation of share equations. Journal of Econometrics, 10(3):361–383, August 1979. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407679900897.

Yancey:1976:MCC

[YJ76] Thomas A. Yancey and George G. Judge. A Monte Carlo comparison of traditional and Stein-rule estimators under squared error loss. Journal of Econometrics, 4(3):285–294, August 1976. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304407676900385.

Zellner:1978:EFP

[Zel78] Arnold Zellner. Estimation of functions of population means and regression coefficients including structural coefficients: A minimum expected loss (MELO) approach. Journal of Econometrics, 8(2): 127–158, October 1978. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407678900246.

Zellner:1974:TSA

[ZP74] Arnold Zellner and Franz Palm. Time series analysis and simultaneous equation econometric models. Journal of Econometrics, 2 (1):17–54, May 1974. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0304407674900281.

Zellner:1973:BAF

[ZW73] Arnold Zellner and Anne D. Williams. Bayesian analysis of the Federal Reserve — MIT–Penn model’s Almon lag consumption function. Journal of Econometrics, 1(3):267–299, October 1973. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407673900109.